### v4

# Joint Denoising and Fusion with Short- and Long-exposure Raw Pairs

Abstract
Denoising and high dynamic range (HDR) imaging are significant yet challenging problems due to the small aperture and sensor size of generic image sensors. Current methods predominantly generate HDR images from a set of bracketed exposure sRGB images. However, they overlook the computational and memory inefficiencies of the Image Signal Processor (ISP) when processing a set of sRGB images with different exposures. Furthermore, the absence of large-scale raw-based HDR datasets limits the research on HDR imaging. Unlike existing methods, the core idea of this work is to utilize the difference between short- and long-exposure images of signal-to-noise ratios to generate HDR images and denoising. To this end, we propose a model tailor-made for double-exposure HDR sensors, leveraging the unique features of the raw data to facilitate raw-to-HDR mapping and raw denoising. Our key insights are threefold: (1) a new computational raw LDR-HDR pair formation pipeline is designed to construct a real-world raw HDR dataset called RealRaw-HDR; (2) a lightweight-efficient HDR model, RepUNet, is developed using the structural reparameterization technique; (3) a plug-and-play alignment-free and motion-aware short-exposure-first selection loss and a colorfulness loss are proposed to mitigate ghost artifacts and color cast. Our empirical evaluation validates the effectiveness of the proposed LDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.
1 Introduction
Figure 1: To address the HDR reconstruction problem, previous methods mainly focus on models trained on processed sRGB datasets, while we promote a novel view to address HDR reconstruction by training on raw images. Furthermore, compared with sRGB-based HDR reconstruction, raw-based HDR reconstruction is more beneficial and reduces the power consumption of ISP, DDR, and NPU by almost half.
Most resource-constrained cameras exhibit low dynamic ranges (LDR), rendering them unable to capture the full range of brightness and color information in real-world scenes. Conversely, high dynamic range imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in LDR images. Despite the availability of dedicated hardware for directly acquiring HDR images, such equipment is typically expensive, thus limiting its practicality for most users. As a result, there has been an increasing focus on fusion-based HDR imaging methods tailored for LDR cameras.
Widely practiced HDR reconstruction methods [16, 5, 27, 37, 39] exploit a set of sRGB images with different exposure levels to produce an HDR image. However, capturing a set of bracketed exposure LDR images can be challenging, especially when the camera is shaking or there are moving objects in the viewfinder frame. [16]. In response to this challenge, Sony has introduced a Digital Overlap (DOL) [41] HDR sensor and a Binned Multiplexed Exposure (BME) HDR sensor [24]. These sensors can shorten the time interval between different exposure images, enabling the “quasi-synchronous” capture of multiple raw images with different exposure times. Nonetheless, the quasi-synchronous mechanism retains relative motion between pre- and post-frames, which can introduce ghost artifacts in the reconstructed HDR results.
Although sRGB-based HDR reconstruction methods [38, 39] have made significant progress in recent years, these methods neglect three critical considerations. Firstly, as shown in Fig. 1, an ISP pipeline needs to be applied separately to different exposure LDR raw images to obtain different exposure LDR sRGB images, which are then reconstructed as an HDR image. The ISP pipeline typically involves complicated processes, including demosaicing, tone mapping, color and contrast tuning, denoising and sharpness, etc. These processes require significant memory and computational resources, which is very unfriendly to resource-limited sensors like security cameras. Secondly, existing approaches [16, 21] substitute the complex camera response curve with a simplistic inverse gamma function to obtain a set of pseudo-linear sRGB images with different exposures. However, this approach may produce non-luminance-aligned images in non-saturated regions during luminance alignment, thus rendering training more challenging. The linear characterization between the different exposure images of the raw images reduces the difficulty of luminance-aligned processing. Thirdly, the raw image data contains more detailed information, and the noise model is more straightforward. Meanwhile, the bit depth of the raw image is generally higher than that of the sRGB image, which is more beneficial to improving the dynamic range of the image. Therefore, the limitations inherent in the current sRGB-based HDR reconstruction methods highlight the necessity of further research and development in the domain of HDR imaging.
Building upon the above observations, we propose reconstructing HDR images from a set of raw LDR images with different exposures (as shown in Fig. 1. Accordingly, we introduce a novel raw LDR-HDR pair formation pipeline and obtain an associated training and testing dataset named RealRaw-HDR. In addition, we devise a simple yet effective raw HDR reconstruction network with reparameterization techniques, establishing a new baseline for the raw HDR reconstruction field. Specifically, upon exploring the HDR sensor imaging system, we discover that simulating the sensor’s Digital Gain can produce the same luminance level as altering the exposure time, given certain conditions. Leveraging this insight, we design a digital imaging pipeline that captures raw images with changing exposure gain based on the camera’s photoelectric signal conversion model. By simulating the HDR sensor imaging system, we built a digital imaging pipeline to obtain raw images with different exposures, thus emerging as a practical tool for generating high-quality LDR-HDR images. The resultant RealRaw-HDR dataset encompasses a substantial collection of LDR-HDR pairs, serving as the foundation for training and evaluation. By incorporating the distinctive attributes of raw images into our approach, we can achieve elevated HDR reconstruction outcomes characterized by enhanced efficiency and accuracy.
To further advance the HDR imaging system, we introduce RepUNet, a lightweight and resource-efficient model for raw HDR reconstruction. RepUNet adopts reparameterization techniques, bypassing the need for computationally intensive alignment modules such as optical flow, deformation convolution, or attention mechanisms, which are commonly employed in existing deep learning-based HDR reconstruction methods [16, 21]. Furthermore, addressing the absence of explicit alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss (AMSS-Loss), which guides the network to focus on local motion regions, effectively mitigating ghost artifacts between short- and long-exposure images. This approach significantly reduces hardware expenses and enhances the real-time performance of HDR imaging systems.
Our contributions are summarized as follows:
  • $\bullet$
    We propose a pioneering pipeline for forming LDR-HDR pairs based on sensor imaging theory, resulting in a large-scale, high-quality raw HDR dataset coined RealRaw-HDR. This dataset serves as the bedrock for follow-up raw HDR reconstruction methodologies.
  • $\bullet$
    We develop a lightweight model, RepUNet, to achieve raw HDR reconstruction, which utilizes structural reparameterization techniques that notably curb the computational demands of hardware devices.
  • $\bullet$
    We introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate ghost artifacts between short- and long-exposure images.
Our contributions signify a substantial advancement in raw HDR image reconstruction research, delivering an efficient and productive solution for producing superior HDR images. Extensive experiments demonstrate that our approach achieves state-of-the-art visual quality and quantitative metrics performance.
2 Related Work
2.1 HDR Imaging
The expansion of camera dynamic range has been a focal point of research over the past decades. One class of technology [33, 31] involves achieving HDR image capture in a single shot by refining digital camera hardware design. The Dual Conversion Gain (DCG) sensor [15], for instance, enhances dynamic range through DCG circuitry, enabling image capture in extremely low-light conditions without compromising performance in high-light scenarios. Sun et al. [33] propose a method for single-exposure HDR imaging through end-to-end optical coding design. Regrettably, these methods often necessitate customization to specific optical systems or image sensors. HDR techniques relying on hardware modification are typically expensive and beyond the reach of average users.
Another class of approaches [25, 21, 36] is time-domain multi-frame fusion HDR, which merges a set of bracketed exposure LDR sRGB images to produce an HDR image. However, this process—capturing images from underexposed to overexposed configurations can result in blurring artifacts for moving objects. Sony has endeavored to address this by introducing the Digital Overlap (DOL) HDR sensor [6]. For instance, dual-exposure DOL sensors like Sony IMX327, IMX385, and OV OS05B produce long-exposure and short-exposure frames in line-interleaving patterns. This pattern significantly shortens the interval between these frames, thus minimizing relative motion. Presently, many methods use DOL sensors to capture a set of bracketed exposure raw images, which ISP processes, and subsequently fusion bracketed exposure LDR sRGB images to reconstruct an HDR image.
Recent trends in HDR imaging involve the application of deep neural networks [18, 26, 2] to learn the production of high-quality HDR images from a set of bracketed exposure LDR images. Kalantari et al. [16] propose a CNN-based HDR approach that employs optical flow to align LDR images before network inference. Wu et al. [36] approach HDR imaging as an image translation problem without explicit motion alignment. Yan et al. [37] introduce spatial attention to achieve LDR image alignment. Liu et al. [21] present an attention-guided deformable convolutional network for multi-frame HDR imaging. Prabhakar et al. [29] introduce an efficient method for generating HDR images using a bilateral guided up-sampler and explore zero-learning for HDR reconstruction. Niu et al. [27] propose a multi-frame HDR imaging method based on generative adversarial learning. Liu et al. [22] propose a Transformer-based [28] HDR imaging method. These deep learning-based approaches [37, 21, 22] consistently elevate state-of-the-art performance. However, these methods reconstruct HDR based on sRGB images at the end of the ISP pipeline and only train on one dataset. They overlook the large computational and storage resources required by the ISP pipeline to process bracketed exposure raw images. It also complicates the ISP system, making it challenging for resource-limited cameras to output high-quality video/images.
2.2 Low-level Raw Image Processing
Learning-based processing of low-level raw images [19, 43] has received increasing attention. Google’s HDR+ [11] produces the raw HDR image by aligning and merging a burst of raw frames with the same low exposure. Zhang et al. [44] construct a real-world super-resolution dataset by designing an optical zoom system and propose a baseline network with a bilateral contextual loss. Qian et al. [30] solve the joint demosaicing, denoising, and super-resolution task with the raw input. Wang et al. [34] propose a lightweight and efficient network for raw image denoising. Sharif et al. [1] offer a new learning-based approach to tackle the challenge of joint demosaicing and denoising on image sensors. Wei et al. [35] investigate the low-light image denoising considering the sensor’s photoelectric properties. Yue et al. [42] achieve state-of-the-art raw image denoising by constructing a dynamic video dataset with noise-clean pairs.
Learning-based raw image processing has demonstrated outstanding potential for high-performance reconstruction from raw sensor data. However, acquiring paired data in the raw domain is difficult and expensive. Our work proposes a new large-scale, high-quality raw dataset and provides a pipeline to acquire raw LDR-HDR pairs based on the imaging system.
3 New raw LDR-HDR pair Formation Pipeline
We begin by analyzing the sensor response of the imaging system and presenting a novel data formation pipeline for raw LDR-HDR pairs based on the camera response model.
3.1 Analysis of CMOS Imaging System
The essence of a CMOS image sensor is photo-electric signal conversion. For a single pixel, the number of electrons $Q$ released during the light-electric signal conversion can be ideally expressed as [12]:
[…]
where $(x,y)$ represents spatial coordinates on the sensor plane, $T$ is the integration time (exposure time), $E(x,y,\lambda)$ signifies the incident spectral irradiance, $S(x,y)$ characterizes the spatial response of the collection site, and $q(\lambda)$ is defined as the ratio (electrons/Joule) of collected electrons per incident light energy for the sensor as a function of wavelength $\lambda$.
Given that $(x,y)$ in Eq. 1 pertains to a single photosensory cell, we assume that each parameter remains constant with respect to position. Consequently, the coordinates $(x,y)$ can be omitted [14]:
[…]
where $\overline{S}$ denotes the expected value of $S(x,y)$ within a single photosensory cell, and $A$ denotes the effective photoreceptor area of the cell.
Subsequently, the camera’s amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [47]:
$D=\frac{K_{\textrm{a}}Q+V_{\textrm{offset}}}{\eta}\times K_{\textrm{d}},$ (3)
where $K_{\textrm{a}}$ represents the Analog Gain, $K_{\textrm{d}}$ stands for the Digital Gain, and $V_{\textrm{offset}}$ accounts for the bias voltage. $\eta$ corresponds to the quantization step associated with the bit depth.
Combining Eq. 2 and Eq. 3, the ideal model for the optical-to-digital conversion is modeled as:
[…]
where $D$ signifies the pixel value in the raw image, $V_{\textrm{offset}}/{\eta}$ accommodates artificially introduced bias voltage to prevent output signals below 0. The raw response value of the bias voltage (i.e., black level) can be directly read out. When the dark current is 00, or we subtract the raw response value of the bias voltage, the pixel value in the raw image can be expressed as:
[…]
We observe from Eq. 5 that under noise-free conditions, adjusting the gain factor ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly adjust the camera’s raw response value. This linear characteristic allows us to achieve an equivalent result as modifying the exposure time ($T$) by simulating the Digital Gain, thereby obtaining a bracketed exposure raw image.
Nonetheless, due to noise in real-world imaging, changing the Digital Gain will also linearly amplify the noise. Therefore, to mitigate noise influence, we avoid this problem as much as possible during data acquisition (in Sec. 3.2.1). Also, we avoid using multiplicative gain to construct long-exposure images that amplify the noise difference between long- and short-exposure images and instead use truncation to similar long-exposure images. In addition, we synthesize the long-short exposure images by measuring the accurate HDR sensor noise model to eliminate the long-short exposure image noise difference. Based on this rationale, we propose an innovative and efficient pipeline for generating LDR-HDR data pairs.
3.2 Formation of Raw LDR-HDR Image Pair
Relative to sRGB images, HDR reconstruction from raw images offers several distinct advantages: it retains a more significant amount of original information and has higher bit depths, which is more beneficial for improving the dynamic range of the image; the Image Signal Processing (ISP) pipeline is more straightforward and less computational demands. These characteristics render raw HDR reconstruction a highly promising approach for deployment in edge devices. Considering these considerations, we create a raw HDR dataset featuring a collection of LDR-HDR data pairs, aptly named RealRaw-HDR.
3.2.1 Data Acquisition
Based on the analysis in Sec. 3.1, we observe that changing the Digital Gain yields a comparable luminance to adjusting the exposure time when the noise level remains sufficiently low. Consequently, we can capture two raw images with identical exposure settings and subsequently manually simulate the Digital Gain to generate two raw images with differing exposures. It’s crucial to emphasize that these two raw images with the same exposure require low noise and no overexposure. In order to capture near-noise-free raw images, we utilize high-end cameras, special photography techniques, and post-processing acquisitions.
Specifically, we use a top-of-the-line FUJI-FILM GFX50S II camera with a wide-aperture lens to capture high-quality raw images. The camera has a $51$ megapixel medium format image sensor with a pixel size of $5.3\mu m$ (The iPhone 14 Pro Max primary camera single pixel size is only $1.22\mu m$). We also set the camera’s ISO to 800 or below and turned on the noise reduction feature to enhance image quality. Subsequently, we subject the raw image to post-processing using the BM3D algorithm, attaining nearly noise-free raw images.
Meanwhile, to simulate the relative motion between long and short exposure images within a real-world scene, we use a human subject to simulate motion between LDR images and trigger the shutter twice in a rapid time interval. To eliminate the risk of unintended camera shake, we securely mount the camera on a tripod and use a remote smartphone to control the shutter release. Following image capture, the raw images undergo black level correction and normalization before being subjected to BM3D processing [23] to effectively reduce noise, thereby yielding raw images nearly devoid of noise.
3.2.2 Data Processing
Fig. 2 shows the proposed data formation pipeline. Here’s an in-depth breakdown of the data formation pipeline:
Selection of exposure time ratio and initial adjustment. Commencing the pipeline, we pack two clean Bayer raw images. We randomly select an exposure time ratio $r$ from {4, 8, 16} at this stage.  [^1] It’s important to note that exposure ratios may not necessarily be multiples of 2. The two normalized raw images, denoted as $I_{1}$ and $I_{2}$, multiply by the maximum pixel value (Max: $2^{12}\times r$). This operation yields $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{2}$ correspondingly.
Long-exposure image simulation ($I_{l}$). Moving forward, we divide $I^{{}^{\prime}}_{1}$ by 1 and then clip the pixel values to a range of 0 to 4095 (12 bit). This operation is equivalent to adjusting the Digital Gain ($K{\textrm{d}}$), achieving an outcome comparable to altering the exposure time. Subsequently, we synthesize the noisy image by long exposure image noise model measured by Sony IMX327 HDR sensor and normalize it. This process simulates a long-exposure noisy image ($I_{l}$) reaching saturation signal level (full well capacity) and the inherent noise generated by the image sensor, which preserves dark detail while losing details in brighter areas.
Short-exposure image simulation ($I_{l}$). Simultaneously, we divide $I^{{}^{\prime}}_{2}$ by $r$ and clip the pixel values to a range of 0 to 4095. (Equivalent to adjusting the Digital Gain). Similar to the previous step, synthesize the short-exposure noisy image by the short-exposure image noise model measured by the Sony IMX327 HDR sensor, and normalization follows suit. This procedure simulates a practical short-exposure image ($I_{s}$). This image retains detail in highlighted portions but loses darker information due to noise interference.
Ground truth image ($I_{gt}$). Additionally, we normalize $I^{{}^{\prime}}_{2}$ to obtain the ground truth image $I_{gt}$. The $I_{gt}$ contains more information on bright regions than $I_{l}$; $I_{gt}$ has a higher signal-to-noise ratio in dark regions compared to $I_{s}$. As a result, our data formation pipeline efficiently generates an extensive array of LDR-HDR data pairs.
Luminance alignment. Finally, after the luminance alignment [16], we obtain the noisy raw LDR images $I^{{}^{\prime}}_{l}$ and $I^{{}^{\prime}}_{s}$, and the corresponding clean raw HDR image $I^{{}^{\prime}}_{gt}$.
The data formation pipeline facilitates the generation of rich LDR-HDR pairs, which is integral to our research and exploration.
Figure 2: The raw LDR-HDR pair formation pipeline. Two clean high-bit raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction, normalization, and BM3D denoising.
Figure 3: Real samples collected by the proposed raw LDR-HDR pair formation pipeline. For display purposes, we do not apply luminance alignment processing.
3.2.3 RealRaw-HDR Dataset
Our dataset is meticulously crafted for dual-exposure HDR sensors, supporting mainstream sensors, including Sony IMX327, IMX385, IMX585, and OV OS05B. To the best of our knowledge, there is an absence of a raw HDR dataset explicitly tailored for these HDR sensors. Furthermore, it is challenging to construct large-scale, high-quality raw HDR datasets from multi-exposed frames. Traditional methods of collecting HDR datasets often lead to misalignment issues, yielding fused HDR images of subpar quality; previous data collection methods necessitate subjects to remain static while capturing multiple exposures, a time-consuming and labor-intensive process. In contrast, our proposed raw LDR-HDR pair formation pipeline is efficient and user-friendly, enabling the effortless creation of a multitude of high-quality data pairs.
Through this pipeline, we gather $240$ pairs of $8192\times 6192$ high-resolution raw LDR-HDR image pairs. Fig. 3 shows an example of two generated LDR-HDR pairs with different exposure ratios. Additionally, by attaching an ISP pipeline to the end of our pipeline, we can create an sRGB-based HDR training dataset. In Tab. 1, we compare the statistics of our dataset with those of other existing HDR datasets.  [^2] If not specified, all raw images in this paper have been processed by a fixed ISP to obtain sRGB versions for visualization purposes.
Table 1: Comparison with existing HDR datasets. ×$\times 3$ indicates that the same set of images can be expanded with three exposure ratios. Theoretically, Our dataset can be expanded $N$ times by the raw LDR-HDR pair formation pipeline, N denotes the number of covered exposure ratios.
Data Quantity Size Format Exposure Ratio
Kalanatri [16] 74 $1490\times 989$ sRGB {4, 8, 16}
Chen [5] 144 $4096\times 2168$ raw, sRGB {4, 8, 16}
Ours $240\times 3$ $8192\times 6192$ raw, sRGB {4, 8, 16}
4 The Proposed RepUNet Network
HDR reconstruction plays a vital role in various applications, such as mobile photography, high-definition displays, and virtual reality, where lightweight and efficient algorithms are highly demanded due to resource limitations. Previous learning-based HDR methods [22, 21, 16] often rely on large and complex models, making them impractical for real-world scenarios. Therefore, we propose a UNet-like lightweight base model, DualUNet, to address this challenge and employ reparameterization techniques to improve its representation capability while balancing storage and computation cost and reconstruction accuracy. We first present the DualUNet and then develop our RepUNet by reparameterization, followed by the design of the loss function.
4.1 Base Model of DualUNet
To ensure high inference speed and cross-device deployment on commodity mobile devices, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a UNet-like network as the base model, namely DualUNet. The overall architecture of the DualUNet is shown in Fig. 4(a). We do not use the computationally demanding explicit alignment in our HDR network for lightweight design. We introduce an alignment-free and motion-aware short-exposure-first selection loss (in Sec. 4.3) that enables training with unaligned pairs to compensate for the absence of alignment modules.
We first use the pixel unshuffle operation [10] to transfer the input raw images $I^{\textquoteright}_{1}$ and $I^{\textquoteright}_{s}$ from $C\times H\times W$ to $4C\times\frac{1}{2}H\times\frac{1}{2}W$ to extract large-scale contextual information while maintaining low computational complexity. Notably, short-exposure images have less motion blur distortion but severe noise in dark areas. Conversely, long-exposure images generally contain more information with a higher signal-to-noise ratio but may suffer from significant motion blur and overexposure.
To better leverage the complementary information from the two types of images, we introduce two distinct sub-encoders: Encoder-S and Encoder-L. Encoder-S extracts features from the short-exposure image, serving as foundational reference features. In parallel, Encoder-L extracts features from the long-exposure image, offering supplementary features. We fuse the multi-scale features using addition or concatenation operations at each feature level to facilitate complementary features to learn motion information from the reference features.
Each Encoder-S consists of a pixel unshuffle $\downarrow 2$ downsampling operation, a $3\times 3$ convolution layer, and a ReLU activation ([Down-Conv-ReLU]). At the same time, the Encoder-L has the structure of [Down-Conv-ReLU-Conv-ReLU]. The pixel unshuffle operation can avoid information loss introduced by conventional downsampling/pooling operations. Finally, the fused features are fed into the decoder to generate the final HDR output.
Refer to caption (a) DualUNet
Refer to caption (b) Topology Convolution Block (TCB)
Figure 4: Illustration of (a) DualUNet and (b) Topology Convolution Block (TCB). In the training phase, the TCB module replaces the normal convolution, which is merged back into a normal 3 × 3 convolution layer in the inference phase.
4.2 RepUNet Network
The representation capability of the plain DualUNet for HDR imaging is significantly limited, as shown in Tab. 6, leading to unsatisfactory performance. To address this concern, we introduce a flexible re-parameterizable module called the Topological Convolution Block (TCB) and replace the normal convolution in DualUNet with the TCB module during the training phase. Notably, the last layer of DualUNet is still a normal convolution.
Although the reparameterization has achieved promising results on other tasks [9, 4, 7, 46], its application in reparameterization blocks for low-level visual tasks remains scarce. We integrate previous reparameterization blocks and flexibly design reparameterization blocks with six operator combinations to enhance the feature representation of the model for HDR and denoising tasks. The TCB consists of several fundamental units: (1) a standard $3\times 3$ convolution for a solid foundation; (2) extending and squeezing convolution to enhance feature expressiveness, (3) Sobel and Laplacian operators for extracting first and second-order spatial derivatives to identify edges; (4) a jump connection to avoid gradient vanishing or exploding, and (5) a $1\times 1$ convolution to encourage information exchange between channels. It’s paramount to underscore that we exclusively employ the TCB with the Laplacian operator within the decoder. This selective approach is grounded in Laplacian operator effectiveness for noise-free images, underpinning its application to enhance feature representation in contexts devoid of noise.
To achieve an efficient HDR network that meets the stipulated design prerequisites of low computational complexity and streamlined hardware device deployment, we simplify the TCB reparameterization into a single $3\times 3$ convolution after training. Following previous works [8, 7, 46], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolution into a normal $3\times 3$ convolution. Additionally, we combine the Sobel and Laplacian operators into a special $3\times 3$ convolution with a fixed convolution kernel. The $1\times 1$ convolution is achieved by padding the convolution kernel with zeros. As a result, TCB can be transformed into a single $3\times 3$ convolution for efficient implementation during the inference stage. By utilizing TCB, we achieve superior HDR results with improved efficiency.
4.3 Loss Functions
Alignment-free and motion-aware short-exposure-first selection loss. In fused-based HDR imaging, eliminating ghosting caused by motion inconsistencies between short and long frames is one of the most challenging issues. Previous work [16, 37] commonly employs optical flow, attention mechanisms, and other methods to establish pixel correspondences between long and short frames. The objective is to suppress ghosting by designing more elaborate fusion strategies. Yet, these motion estimation and alignment methods are often the most computationally intensive components and cannot be accommodated by the current level of hardware design. To ensure fast inference and easy hardware deployment, we abandon explicit motion alignment schemes and instead explore a more efficient alternative. Consequently, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate the ghost artifacts.
Unlike other tasks involving image alignment, such as video motion estimation and stereo matching, long and short-exposure image fusion in HDR reconstruction does not necessarily require pixel-level correspondence in real-world scenes. The reason is that it is challenging to recover sharp object edges due to motion blur. In contrast, short-exposure images exhibit less motion blur distortion. Therefore, ghost artifacts can be suppressed by simply detecting motion regions in long-frame images through some mechanism, discarding these pixels during the fusion process, and relying solely on the corresponding regions in short-exposure images as the exclusive information source for fusion. Based on the same consideration, overexposed regions in the long-exposure image should likewise be discarded from the fusion process. Therefore, in practical engineering applications, short exposures are often used as reference images.
Based on the above analysis, we devise the following strategies for double-exposure HDR fusion: in scenarios involving motion or overexposure within the fused region, we prefer to select the short-exposure image that contains more information; when the signal-to-noise ratio of the short frame is too low, we prefer to select the long-exposure image that contains more information. It is worth noting that our strategy is inclined to address motion artifacts with a higher priority than low signal-to-noise ratios. Consequently, when both ghost artifacts and low signal-to-noise ratios are concurrently evident, the short-exposure image is chosen to prioritize the motion artifacts.
Figure 5: An illustrative sample of data construction for the proposed pseudo location-based motion alignment loss.
Figure 6: Visual comparison of state-of-the-art HDR reconstruction methods on our synthetic test dataset.
Note that after the luminance alignment, the long and short frames already have similar luminance levels, and the regions with remarkable luminance differences are mainly caused by motion and overexposure. Therefore, we expect the network to perceive the motion and overexposed regions by focusing on the luminance gap between the long and short exposures, thereby realizing the short-frame prioritization strategy mentioned above. For this purpose, we design an alignment-free and motion-aware short-exposure-first selection loss function.
$L_{\textrm{AMSS}}=1-MSSSIM(\widetilde{I}^{out}\odot M,{I}^{gt}\odot M),$ (6)
where $\odot$ denotes the point-wise multiplication, and $M$ is a binary mask with “1” for motion regions in long frames, and “0” otherwise. Given a mask $M$ indicating motion and overexposed regions, the above loss formula implements a strategy that encourages short-frame prioritization.
To avoid explicit motion alignment, we adopt the synthetic motion to construct the mask $M$ in the training quadruplets {$I^{{}^{\prime}}_{l}$, $I^{{}^{\prime}}_{s}$, $M$, $I^{{}^{\prime}}_{gt}$. Specifically, as shown in Fig. 5, for each patch of an image, we randomly select a rectangle (40 to 60-pixel range) of random length and width from the long-exposure image and then move and overlap it to a random location in the range of -30 to 30 relative to the patch. The patch regions before and after the movement are labeled as 1s in the mask $M$. By introducing the mask $M$ to guide the network to focus on moving and overexposed regions, the model effectively prioritizes the short-exposure information over the long-exposed counterpart within these regions. The masks are used only in training, and the inference stage inputs only short and long exposure frames.
Reconstruction loss. For saturated areas, the L2 loss punishes any deviation of pixel values from the ground truth. This allows the model to select short-exposure information in over-exposed areas. To achieve the best HDR reconstruction results, we employ the MS-SSIM loss guide model, which learns short-exposure image information for global motion. By combining these loss functions, our model effectively produces superior results for both areas with motion and saturated regions.
$L_{\textrm{pix}}=\|\widetilde{I}^{out}-{I}^{gt}\|_{2},$ (7)
$L_{\textrm{ssim}}=1-MSSSIM(\widetilde{I}^{out},{I}^{gt}).$ (8)
Table 2: Performance comparison of different HDR models on our synthetic dataset. Params and FLOPs represent the total number of network parameters and floating-point operations. The FLOPs and Run Times results are measured on an RTX 3090 device with a resolution of $4096\times 2952$ raw images. Metrics with $\uparrow$ and $\downarrow$ denote higher better and lower better, respectively. The best and second-best performances are in bold and underlined, respectively. "-" indicates the result is not available.
Methods FLOPs Params Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]
DeepHDR[16] 2409.32G 15.26M 4.3 ms 43.3680 1.3767 43.5551 1.3844 43.7312 1.3679 42.8178 1.3779
NHDRRNet[40] 826.17G 40.26M 7.9 ms 33.0206 2.7308 33.0127 2.7277 33.0392 2.7203 33.0101 2.7443
UNet-SID[3] 640.89G 7.76M 3.1 ms 44.3892 1.3434 44.3314 1.3535 44.4551 1.3312 44.3811 1.3456
SGN[10] 712.66G 4.78M 3.3 ms 44.6094 1.3235 44.5074 1.3398 44.7078 1.3067 44.6131 1.3240
HDR-Transformer[22] 3698.28G 1.23M - 44.3895 1.3681 44.3311 1.3811 44.4140 1.3619 44.4235 1.3613
AHDRNet[37] 2848.29G 0.93M 23.6 ms 44.7985 1.2939 44.8548 1.2957 44.8343 1.2892 44.7064 1.2968
Ours 127.55G 0.82M 2.9 ms 44.8081 1.2886 44.7575 1.3000 44.8482 1.2812 44.8187 1.2842
Figure 7: Visual comparisons with the state-of-the-art methods on HDR sensor raw dataset.
Bayer loss. We propose a color correction loss, named Bayer loss, to minimize color cast and artifacts. We average the two G channels of the output (RGGB pattern) and ground truth (RGGB pattern) respectively, and then concatenate the averaged G channel with the R and B channels to perform a naive transformation to the RGB color space, producing two RGB images: $\widetilde{I}^{out}_{rgb}$ and ${I}^{gt}_{rgb}$. Then, we impose the colorfulness loss between the processed output and ground truth by the cosine embedding loss.
$L_{\textrm{b}}=\textrm{Cosine}(\widetilde{I}_{rgb}^{out},{I}_{rgb}^{gt}),$ (9)
where Cosine denotes cosine embedding loss [32]. The overall loss function is
$L=\alpha\cdot L_{\textrm{AMSS}}+\beta\cdot L_{\textrm{b}}+\gamma\cdot L_{\textrm{pix}}+\eta\cdot L_{\textrm{ssim}}.$ (10)
where $\alpha$, $\beta$, $\gamma$, and $\eta$ are the corresponding weight coefficients.
5 Experiments
5.1 Experimental Setup
Datasets and Metrics. We utilize the proposed RealRaw-HDR dataset for training. For testing, we first evaluate our method on the dataset, which is synthesized using the data formation pipeline. This dataset was not utilized during training. This test set contains 30 sets of test samples containing different exposure ratios (i.e., 4, 8, and 16) with a resolution of $4096\times 2176$. To validate the generalization ability of our method on real data, we utilize a FUJI-FILM GFX50S II camera to capture seven sets of real-world bracketed exposure raw images and the corresponding static images for generating the ground truth. Furthermore, to validate the generalization ability of our method on HDR sensors, we utilize Chen’s test dataset [5] for cross-validation, which has short-long raw images captured by a Sony IMX267 image sensor.
We perform a quantitative evaluation using the PSNR, SSIM, and CIE L*a*b* space  [^3] CIE L*a*b* is a color space specified by the International Commission on Illumination. It describes all the colors visible to the human eye and was created to serve as a device-independent model to be used as a reference. [45, 13, 20] (also known as $\Delta E$). $\Delta E$ can effectively assess chromaticity, contrast, and color accuracy variations within HDR images rather than exclusively concentrating on luminance differences (HDR-VDP-2). It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE Lab color space.
$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (11)
where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE Lab version of the predicted HDR image and ground truth, respectively.
Table 3: Performance comparison of different HDR models on the actual HDR sensor raw dataset [5]. The best and second-best performances are in bold and underlined, respectively.
Methods All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]
DeepHDR[16] 39.4902 0.9731 2.0670 39.2987 0.9716 2.1201 40.3268 0.9779 1.9159 38.8450 0.9697 2.1648
NHDRRNet[40] 30.4292 0.9640 5.2132 30.5489 0.9615 5.0771 30.6833 0.9704 5.1679 30.0553 0.9601 5.3945
UNet-SID[3] 39.6099 0.9735 2.1527 39.4473 0.9723 2.1860 40.4429 0.9772 1.9640 38.9394 0.9712 2.3081
SGN[10] 39.3674 0.9727 2.3317 39.3531 0.9718 2.3357 40.0126 0.9761 2.1956 38.7366 0.9704 2.4639
HDR-Transformer[22] 39.9483 0.9726 2.1241 39.7823 0.9715 2.1859 40.5929 0.9750 2.0068 39.4698 0.9713 2.1793
AHDRNet[37] 40.4131 0.9695 2.0123 40.4692 0.9677 1.9829 41.0748 0.9717 1.8519 39.6953 0.9692 2.2025
Ours 40.5238 0.9747 1.9568 40.4061 0.9733 1.9743 41.4010 0.9788 1.7974 39.7642 0.9721 2.0988
Figure 8: Visual comparison of state-of-the-art HDR reconstruction methods on RealRaw-HDR dataset with realistic exposure ratios.
Implementation Details. We train our model using the Adam optimizer [17] with weight decay $1\times 10^{4}$, learning rate $10^{-4}$, and $\beta_{1}$ and $\beta_{2}$ values set to 0.9 and 0.999, respectively. The input patch size for the network is $256\times 256$, and the batch size is 32. Our model is implemented in PyTorch and trained with an NVIDIA RTX 3090 GPU.
5.2 Comparison with the State-of-the-art Methods
We choose several representative low-level vision methods for comparisons, including four HDR methods based on sRGB images (AHDRNet [37], DeepHDR [16], NHDRRNet [40], HDR-Transformer[22]), as well as two methods for denoising raw images (SGN [10] and UNet-SID [3]). For fair comparisons, we re-train all the methods using the RealRaw-HDR dataset. Additionally, for AHDRNet, DeepHDR, HDR-Transformer, and NHDRRNet, we modify the network inputs to accommodate dual-exposure raw images. Similarly, for SGN and UNet-SID, we concatenate the long- and short-exposure images as inputs.
Table 4: Quantitative comparisons of different loss functions. AMSS-Loss represents the alignment-free and motion-aware short-exposure-first selection loss.
ID Method Bayer-Loss AMSS-Loss All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]
1 RepUNet 39.6403 2.1479 39.5214 2.1577 40.5216 1.9603 38.8780 2.3256
2 RepUNet 40.0251 2.1162 39.9900 2.1143 40.8125 1.9700 39.2729 2.2642
3 RepUNet 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
Evaluation on synthetic dataset. We first evaluated our method on a synthetic dataset generated using the raw HDR data formation pipeline. The quantitative comparison results are shown in Tab. 2. The results clearly show that our method outperforms previous methods in almost all metrics on the synthetic dataset. Notably, our lightweight-efficient RepUNet model has fewer parameters (0.82M) and remarkably low GFLOPs (127G FLOPs). This efficiency allows us to process two 4K Bayer raw images at only 2.9ms using an NVIDIA RTX 3090 GPU. In contrast, other models with comparable performance necessitate significantly longer processing times. RepUNet achieves comparable performance with AHDRNet with only $4.5\%$ of its computational complexity (127G vs. 2848G). Fig 6 visualizes that our method can effectively eliminate noise and ghosting artifacts in the reconstructed HDR. In comparison, DeepHDR [16], NHDRRNet [40], and SGN [10] exhibit numerous artifacts in the palm motion region. However, our proposed RepUNet can reconstruct HDR images without ghosting (see rows 2 and 4).
Evaluation on HDR sensor raw dataset. To validate the generalization ability of our method on the real-world HDR sensor, we utilize Chen’s test dataset [5] for cross-validation, which has raw images captured by a Sony IMX267 image sensor. Compared with previous methods, our method achieves state-of-the-art performance in terms of both visual quality and quantitative metrics. The visual results from tests on the HDR sensor raw dataset (as shown in Fig. 7) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from severe color casts. Furthermore, results in Tab. 3 reveal that compared to AHDRNet [37], our method yields an improvement of more than 0.35 dB and 0.05 gain in PSNR and $\Delta E$, respectively, for scenes with an exposure ratio of 8. On average, our method attains gains exceeding 0.1 dB and 0.15 in these metrics.
Evaluation on FUJI raw dataset. We then evaluate our method on the FUJI raw datasets, which are real-world bracketed exposure raw images captured by the FUJI-FILM GFX50S II camera. Fig. 8 compares results from two high dynamic range scenes, where our method achieves significantly better visualization. Our method can recover both fine details in overexposed regions and rich colors in underexposed areas without introducing artifacts (see rows 1 and 2). Compared to AHDRNet, our method can effectively remove noise and preserve the structure of dark regions. Notably, the alignment module in DeepHDR, AHDRNet, HDR-Transformer, and NHDRRNet requires many line buffers, making it challenging to deploy on resource-limited edge devices. Notably, HDR-Transformer fails to perform inference even on RTX 3090 devices. In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images, as evident in the visual results of Fig. 6, Fig. 7, and Fig. 8.
Table 5: Performance comparison of different HDR models on RealRaw-HDR dataset with realistic exposure ratios. The FLOPs are measured on the raw image of $7808\times 5824$ resolution. The best and second-best performances are in bold and underlined, respectively.
[…]
DeepHDR[16] 8987.98G 40.3610 0.9740 1.5235
NHDRRNet[40] 3081.94G 33.3725 0.9694 2.8789
UNet-SID[3] 2390.83G 41.9495 0.9751 1.4881
SGN[10] 2658.57G 41.8942 0.9750 1.4874
HDR-Transformer[22] 13946.42G 41.9963 0.9753 1.4897
AHDRNet[37] 10625.58G 42.6409 0.9763 1.4004
Ours 475.83G 42.5364 0.9760 1.4101
Table 6: Reparameterization ablation results. The FLOPs and Run Times are measured on the raw image of 4K resolution.
Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]
DualUNet 93.26G 0.82M 3.0 ms 39.7941 2.1202 39.8026 2.1036 40.5092 1.9790 39.0705 2.2780
[…]
RepUNet 93.26G 0.82M 2.9 ms 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
Despite the fact that all models are trained on the RealRaw-HDR dataset, which is synthesized using the data formation pipeline, they consistently excel on both the synthetic test set and the real-world dataset. Particularly noteworthy is the remarkable performance achieved on the test set comprised of raw images captured by the HDR sensor [5]. These results serve as solid evidence of the generalizability of our proposed RealRaw-HDR dataset and the HDR data formation pipeline. It underscores the robustness inherent in our approach and its capacity to tackle real-world challenges. This not only validates the efficacy of our methodology but also highlights its considerable potential for diverse real-world applications.
5.3 Ablation Study
This section investigates the raw LDR-HDR pair formation pipeline and the importance of different components in the whole RepUNet. We ablate the baseline model step by step and compare the performance differences.
Generalization of our LDR-HDR pair formation pipeline. Our raw LDR-HDR pair formation pipeline is proposed to generate paired raw LDR-HDR data but can also be adapted to generate paired sRGB HDR data. To demonstrate such generalization, we transform the collected RealRaw-HDR dataset with a fixed ISP pipeline into the sRGB color space, named the Raw2RGB-HDR dataset. For comparison, we train the sRGB HDR method AHDRNet [37] on our Raw2RGB-HDR dataset (240 pairs) and Kalantari’s dataset [16] (taking the first two exposures as input, 74 pairs of images), respectively. The test dataset is from Kalantari’s dataset. Results in Tab. 7 show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on Kalantari’s dataset by 2.86 dB in PSNR. The performance gains benefit from an efficient and user-friendly data acquisition pipeline that generates more trainable data pairs. The results demonstrate that our data pipeline is also effective in generating paired LDR-HDR data in sRGB space.
Table 7: We train the sRGB HDR method AHDRNet on our Raw2RGB-HDR dataset (240 pairs) and Kalantari’s dataset, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR dataset.
Method Dataset PSNR PSNR-$\mu$
AHDRNet Kalantari 35.4581 38.1618
Raw2RGB-HDR 38.3183 39.8896
Loss Functions. To test the effects of alignment-free and motion-aware short-exposure-first selection loss and Bayer loss, we set the $L2$ joint $L_{\textrm{ssim}}$ loss as the baseline loss and step-by-step modify the loss function combination. Tab. 4 and 9 show that adding the AMSS loss and Bayer loss steadily improves visual quality and quantitative results. RepUNet with joint loss achieves the best results, outperforming the baseline by 0.5 dB in PSNR and by 0.16 in $\Delta E$ on average. As Fig. 9 shows, alignment-free and motion-aware short-exposure-first selection loss (AMSS-Loss) effectively suppresses the ghosting artifacts (see columns 3 and 4). Meanwhile, our proposed Bayer loss can alleviate the color cast (see columns 2 and 3).
Figure 9: Visual results of RepUNet and its baseline variants. By combining these loss functions, our model can effectively produce top-notch results for both motion and saturated areas.
[…]
6 Discussion and Future Work
Although our method can produce high-quality HDR images in real scenes, it has several limitations that may inspire future work.
Our data synthesis pipeline employs noise simulation to synthesize long and short-exposure images. However, we have found that if the noise distribution of the images we test is significantly different from the noise model we have measured, the results of the reconstruction are often unsatisfactory. Nevertheless, for practical engineering applications, it is important to accurately measure each type of image sensor noise model to achieve efficient raw denoising.
Furthermore, our work does not cover scenarios involving flicker. This is because the flicker problem generally arises in short-exposure images, and it is difficult to determine which region is better when flicker and motion occur simultaneously. Unfortunately, there is no clear-cut solution that can effectively address scenes where both flicker and motion are present. To overcome this challenge, it may be necessary to create a flicker model to generate a substantial amount of training data.
7 Conclusion
We propose a novel data formation pipeline for raw LDR-HDR pairs based on HDR sensor imaging systems and construct a new high-quality raw dataset and raw LDR-HDR pairs dataset. This dataset can efficiently support many kinds of dual-exposure HDR sensor development, which provides vital data support for realizing the new generation of AI ISPs. Meanwhile, we develop a lightweight and efficient RepUNet network that uses structural reparameterization for HDR image reconstruction. Our evaluation has shown that this model can be easily deployed on mobile devices. We propose a plug-and-play alignment-free and motion-aware short-exposure-first selection loss and Bayer loss to address the issue of ghosting artifacts and color cast. Extensive experiments demonstrate that our approach achieves state-of-the-art performance in both visual quality and quantitative metrics.

### v5

@@
-Joint Denoising and Fusion with Short- and Long-exposure Raw Pairs
+Efficient HDR Reconstruction from Real-World Raw Images
 Abstract
-Denoising and high dynamic range (HDR) imaging are significant yet challenging problems due to the small aperture and sensor size of generic image sensors.
-Current methods predominantly generate HDR images from a set of bracketed exposure sRGB images.
-However, they overlook the computational and memory inefficiencies of the Image Signal Processor (ISP) when processing a set of sRGB images with different exposures.
-Furthermore, the absence of large-scale raw-based HDR datasets limits the research on HDR imaging.
-Unlike existing methods, the core idea of this work is to utilize the difference between short- and long-exposure images of signal-to-noise ratios to generate HDR images and denoising.
-To this end, we propose a model tailor-made for double-exposure HDR sensors, leveraging the unique features of the raw data to facilitate raw-to-HDR mapping and raw denoising.
-Our key insights are threefold: (1) a new computational raw LDR-HDR pair formation pipeline is designed to construct a real-world raw HDR dataset called RealRaw-HDR; (2) a lightweight-efficient HDR model, RepUNet, is developed using the structural reparameterization technique; (3) a plug-and-play alignment-free and motion-aware short-exposure-first selection loss and a colorfulness loss are proposed to mitigate ghost artifacts and color cast.
-Our empirical evaluation validates the effectiveness of the proposed LDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.
+The widespread usage of high-definition screens on edge devices stimulates a strong demand for efficient high dynamic range (HDR) algorithms.
+However, many existing HDR methods either deliver unsatisfactory results or consume too much computational and memory resources, hindering their application to high-resolution images (usually with more than 12 megapixels) in practice.
+In addition, existing HDR dataset collection methods often are labor-intensive.
+In this work, in a new aspect, we discover an excellent opportunity for HDR reconstructing directly from raw images and investigating novel neural network structures that benefit the deployment of mobile devices.
+Our key insights are threefold: (1) we develop a lightweight-efficient HDR model, RepUNet, using the structural re-parameterization technique to achieve fast and robust HDR; (2) we design a new computational raw HDR data formation pipeline and construct a real-world raw HDR dataset, RealRaw-HDR; (3) we propose a plug-and-play motion alignment loss to mitigate motion ghosting under limited bandwidth conditions.
+Our model contains less than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU.
+While being highly efficient, our model also outperforms the state-of-the-art HDR methods in terms of PSNR, SSIM, and a color difference metric.
 1 Introduction
-Figure 1: To address the HDR reconstruction problem, previous methods mainly focus on models trained on processed sRGB datasets, while we promote a novel view to address HDR reconstruction by training on raw images.
-Furthermore, compared with sRGB-based HDR reconstruction, raw-based HDR reconstruction is more beneficial and reduces the power consumption of ISP, DDR, and NPU by almost half.
 Most resource-constrained cameras exhibit low dynamic ranges (LDR), rendering them unable to capture the full range of brightness and color information in real-world scenes.
-Conversely, high dynamic range imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in LDR images.
+Conversely, high dynamic range (HDR) imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in LDR images.
 Despite the availability of dedicated hardware for directly acquiring HDR images, such equipment is typically expensive, thus limiting its practicality for most users.
-As a result, there has been an increasing focus on fusion-based HDR imaging methods tailored for LDR cameras.
-Widely practiced HDR reconstruction methods [16, 5, 27, 37, 39] exploit a set of sRGB images with different exposure levels to produce an HDR image.
-However, capturing a set of bracketed exposure LDR images can be challenging, especially when the camera is shaking or there are moving objects in the viewfinder frame.
-[16].
-In response to this challenge, Sony has introduced a Digital Overlap (DOL) [41] HDR sensor and a Binned Multiplexed Exposure (BME) HDR sensor [24].
-These sensors can shorten the time interval between different exposure images, enabling the “quasi-synchronous” capture of multiple raw images with different exposure times.
-Nonetheless, the quasi-synchronous mechanism retains relative motion between pre- and post-frames, which can introduce ghost artifacts in the reconstructed HDR results.
-Although sRGB-based HDR reconstruction methods [38, 39] have made significant progress in recent years, these methods neglect three critical considerations.
-Firstly, as shown in Fig.
-1, an ISP pipeline needs to be applied separately to different exposure LDR raw images to obtain different exposure LDR sRGB images, which are then reconstructed as an HDR image.
-The ISP pipeline typically involves complicated processes, including demosaicing, tone mapping, color and contrast tuning, denoising and sharpness, etc.
-These processes require significant memory and computational resources, which is very unfriendly to resource-limited sensors like security cameras.
-Secondly, existing approaches [16, 21] substitute the complex camera response curve with a simplistic inverse gamma function to obtain a set of pseudo-linear sRGB images with different exposures.
-However, this approach may produce non-luminance-aligned images in non-saturated regions during luminance alignment, thus rendering training more challenging.
-The linear characterization between the different exposure images of the raw images reduces the difficulty of luminance-aligned processing.
-Thirdly, the raw image data contains more detailed information, and the noise model is more straightforward.
-Meanwhile, the bit depth of the raw image is generally higher than that of the sRGB image, which is more beneficial to improving the dynamic range of the image.
-Therefore, the limitations inherent in the current sRGB-based HDR reconstruction methods highlight the necessity of further research and development in the domain of HDR imaging.
-Building upon the above observations, we propose reconstructing HDR images from a set of raw LDR images with different exposures (as shown in Fig.
-1.
-Accordingly, we introduce a novel raw LDR-HDR pair formation pipeline and obtain an associated training and testing dataset named RealRaw-HDR.
-In addition, we devise a simple yet effective raw HDR reconstruction network with reparameterization techniques, establishing a new baseline for the raw HDR reconstruction field.
-Specifically, upon exploring the HDR sensor imaging system, we discover that simulating the sensor’s Digital Gain can produce the same luminance level as altering the exposure time, given certain conditions.
-Leveraging this insight, we design a digital imaging pipeline that captures raw images with changing exposure gain based on the camera’s photoelectric signal conversion model.
-By simulating the HDR sensor imaging system, we built a digital imaging pipeline to obtain raw images with different exposures, thus emerging as a practical tool for generating high-quality LDR-HDR images.
-The resultant RealRaw-HDR dataset encompasses a substantial collection of LDR-HDR pairs, serving as the foundation for training and evaluation.
-By incorporating the distinctive attributes of raw images into our approach, we can achieve elevated HDR reconstruction outcomes characterized by enhanced efficiency and accuracy.
-To further advance the HDR imaging system, we introduce RepUNet, a lightweight and resource-efficient model for raw HDR reconstruction.
-RepUNet adopts reparameterization techniques, bypassing the need for computationally intensive alignment modules such as optical flow, deformation convolution, or attention mechanisms, which are commonly employed in existing deep learning-based HDR reconstruction methods [16, 21].
-Furthermore, addressing the absence of explicit alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss (AMSS-Loss), which guides the network to focus on local motion regions, effectively mitigating ghost artifacts between short- and long-exposure images.
-This approach significantly reduces hardware expenses and enhances the real-time performance of HDR imaging systems.
+As a result, there has been an increasing focus on fusion-based HDR imaging methods.
+Recent methods [35, 20, 4, 25, 34] based on convolution neural networks (CNNs) [17, 22] have made impressive progress in HDR reconstruction performance, thanks to their scalability and flexibility from constructing elementary building blocks like convolutional layers.
+However, superior performance is usually obtained at a cost of heavy computational burden [20, 34, 21].
+Although this can be alleviated by elaborate network structures or dedicated computing engines (e.g., GPU and NPU), the hardware cost and power consumption still limit the deployment of existing deep HDR reconstruction networks.
+Specifically, the growing number of high-definition screens on edge devices (e.g., smartphones, security cameras, and televisions) calls for a practical HDR reconstruction solution.
+On the other hand, in the image processing pipeline, HDR reconstruction is widely used in the sRGB domain.
+Previous methods [35, 22, 36] exploit a set of sRGB images with different exposure levels to produce an HDR image, which has made rapid development in recent years.
+However, they tend to overlook three critical aspects.
+1) Dataset Collection: Existing methods [3, 46] follow Kalantari et al.
+[14] to construct datasets.
+They first make the subject static and take three sets of images with different exposures, and then make the subject move twice to take dynamic images with different exposures.
+However, this process is labor intensive and difficult to acquire on a large scale.
+2) ISP Processing Speed: When obtaining raw LDR images with different exposures, the ISP pipeline must be performed separately on each exposure.
+This incurs additional memory and computational overhead and leads to lower frame rates for HDR image output.
+3) Reconstruction Quality: Raw images contain more delicate details of the original sensor signal that can be lost while processing sRGB images.
+The limitations of current HDR reconstruction methods highlight the need for further research and development in HDR imaging.
+In this paper, we propose an efficient scheme for HDR image reconstruction in the raw image domain.
+By analyzing the HDR image sensor system, we design a lightweight and efficient model for raw HDR reconstruction named RepUNet.
+RepUNet adopts reparameterization techniques and does not contain explicit computationally expensive alignment modules, such as optical flow [14], deformation convolution [4], or attention [20, 37], which are commonly used in existing deep learning-based HDR reconstruction methods [14, 20, 33, 24].
+To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss, which encourages the network to focus on local motion patterns and alleviate misalignment between short- and long-exposure images.
+Consequently, our approach significantly reduces hardware costs and improves the real-time performance of HDR imaging systems.
+To further promote the HDR imaging system, we investigate the HDR sensor imaging principle.
+We observe that changing the Gain of the image sensor can have a similar effect as modifying the exposure time under noise-free conditions.
+Leveraging this insight, we design an automatic control imaging system that captures raw images with different exposures, based on a digital camera photoelectric signal conversion model.
+This automatic control operable system satisfies real-world scenes’ dynamic range requirements, making it a practical tool for generating high-quality HDR images.
+The resulting RealRaw-HDR dataset includes many LDR-HDR pairs for training and evaluation.
+By incorporating the unique characteristics of raw images into our approach, we can achieve superior HDR reconstruction results with increased efficiency and accuracy.
 Our contributions are summarized as follows:
   • $\bullet$
-    We propose a pioneering pipeline for forming LDR-HDR pairs based on sensor imaging theory, resulting in a large-scale, high-quality raw HDR dataset coined RealRaw-HDR.
-This dataset serves as the bedrock for follow-up raw HDR reconstruction methodologies.
+    We investigate the structure re-parameterizable technique for the HDR task and propose a lightweight model, RepUNet, with Topological Convolution Block (TCB).
+TCB can be used to improve the HDR performance of any HDR model without introducing any extra burden for inference.
   • $\bullet$
-    We develop a lightweight model, RepUNet, to achieve raw HDR reconstruction, which utilizes structural reparameterization techniques that notably curb the computational demands of hardware devices.
+    We introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate ghost artifacts.
   • $\bullet$
-    We introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate ghost artifacts between short- and long-exposure images.
-Our contributions signify a substantial advancement in raw HDR image reconstruction research, delivering an efficient and productive solution for producing superior HDR images.
-Extensive experiments demonstrate that our approach achieves state-of-the-art visual quality and quantitative metrics performance.
+    We propose a novel computational photography-based pipeline for raw HDR image formation and construct a real-world raw HDR dataset, i.e., RealRaw-HDR.
+Our contributions represent a significant step forward in raw HDR image reconstruction research, providing an effective and efficient solution for producing high-quality HDR images.
+Our model contains less than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU.
+While highly efficient, our model also outperforms the state-of-the-art HDR methods by a large margin in terms of PSNR, SSIM, and a color difference metric.
 2 Related Work
 2.1 HDR Imaging
-The expansion of camera dynamic range has been a focal point of research over the past decades.
-One class of technology [33, 31] involves achieving HDR image capture in a single shot by refining digital camera hardware design.
-The Dual Conversion Gain (DCG) sensor [15], for instance, enhances dynamic range through DCG circuitry, enabling image capture in extremely low-light conditions without compromising performance in high-light scenarios.
-Sun et al.
-[33] propose a method for single-exposure HDR imaging through end-to-end optical coding design.
-Regrettably, these methods often necessitate customization to specific optical systems or image sensors.
-HDR techniques relying on hardware modification are typically expensive and beyond the reach of average users.
-Another class of approaches [25, 21, 36] is time-domain multi-frame fusion HDR, which merges a set of bracketed exposure LDR sRGB images to produce an HDR image.
-However, this process—capturing images from underexposed to overexposed configurations can result in blurring artifacts for moving objects.
-Sony has endeavored to address this by introducing the Digital Overlap (DOL) HDR sensor [6].
-For instance, dual-exposure DOL sensors like Sony IMX327, IMX385, and OV OS05B produce long-exposure and short-exposure frames in line-interleaving patterns.
-This pattern significantly shortens the interval between these frames, thus minimizing relative motion.
-Presently, many methods use DOL sensors to capture a set of bracketed exposure raw images, which ISP processes, and subsequently fusion bracketed exposure LDR sRGB images to reconstruct an HDR image.
-Recent trends in HDR imaging involve the application of deep neural networks [18, 26, 2] to learn the production of high-quality HDR images from a set of bracketed exposure LDR images.
+Recently, benefiting from the fast development of deep learning techniques, training deep neural networks for effective HDR reconstruction has become increasingly popular.
+Many methods apply deep neural networks [16, 24, 29, 13] to learn the production of high-quality HDR images from a set of bracketed exposure LDR images.
 Kalantari et al.
-[16] propose a CNN-based HDR approach that employs optical flow to align LDR images before network inference.
+[14] proposed a CNN-based HDR approach that employs optical flow to align LDR sRGB images before network inference.
 Wu et al.
-[36] approach HDR imaging as an image translation problem without explicit motion alignment.
+[33] approached HDR imaging as an image translation problem without explicit motion alignment.
 Yan et al.
-[37] introduce spatial attention to achieve LDR image alignment.
+[34] introduced spatial attention to achieve LDR image alignment.
 Liu et al.
-[21] present an attention-guided deformable convolutional network for multi-frame HDR imaging.
+[20] presented an attention-guided deformable convolutional network for multi-frame HDR imaging.
 Prabhakar et al.
-[29] introduce an efficient method for generating HDR images using a bilateral guided up-sampler and explore zero-learning for HDR reconstruction.
+[27] introduced an efficient method for generating HDR images using a bilateral guided up-sampler and exploring zero-learning for HDR reconstruction.
 Niu et al.
-[27] propose a multi-frame HDR imaging method based on generative adversarial learning.
+[25] proposed a multi-frame HDR imaging method based on generative adversarial learning.
 Liu et al.
-[22] propose a Transformer-based [28] HDR imaging method.
-These deep learning-based approaches [37, 21, 22] consistently elevate state-of-the-art performance.
+[21] proposed a Transformer-based [26] HDR imaging method.
+These deep learning-based approaches [34, 20, 21] consistently elevated state-of-the-art performance.
 However, these methods reconstruct HDR based on sRGB images at the end of the ISP pipeline and only train on one dataset.
-They overlook the large computational and storage resources required by the ISP pipeline to process bracketed exposure raw images.
+They overlook the large computational and storage resources the ISP pipeline requires to process bracketed exposure raw images.
 It also complicates the ISP system, making it challenging for resource-limited cameras to output high-quality video/images.
+To simplify the ISP system, another class of HDR reconstruction methods is based on raw image input.
+Google HDR+ produced the raw HDR image by aligning and merging a burst of raw frames with the same low exposure.
+Nevertheless, this approach requires a complex ISP system design and takes up a lot of DDR memory.
+Zou et al.
+[46] proposed reconstructing HDR images from a single raw image and collecting a raw/HDR paired dataset.
+However, this dataset is not suitable for real HDR sensors.
+Therefore, none of the existing methods can meet the requirements of real scenarios.
 2.2 Low-level Raw Image Processing
-Learning-based processing of low-level raw images [19, 43] has received increasing attention.
-Google’s HDR+ [11] produces the raw HDR image by aligning and merging a burst of raw frames with the same low exposure.
+Due to the merits of raw data, raw-based image processing [18, 41] has made significant progress in recent years.
+The work in [40] first performs the demoireing task in the raw domain and then utilizes a pretrained ISP module to transform the result into the sRGB domain.
+Yang et al.
+[38] proposed a single-stage network empowered by feature domain adaptation to decouple the denoising and color mapping tasks in raw low-light enhancement.
 Zhang et al.
-[44] construct a real-world super-resolution dataset by designing an optical zoom system and propose a baseline network with a bilateral contextual loss.
+[42] constructed a real-world super-resolution dataset by designing an optical zoom system and proposed a baseline network with a bilateral contextual loss.
 Qian et al.
-[30] solve the joint demosaicing, denoising, and super-resolution task with the raw input.
+[28] solved the joint demosaicing, denoising, and super-resolution task with the raw input.
 Wang et al.
-[34] propose a lightweight and efficient network for raw image denoising.
+[30] proposed a lightweight and efficient network for raw image denoising.
 Sharif et al.
-[1] offer a new learning-based approach to tackle the challenge of joint demosaicing and denoising on image sensors.
+[1] proposed a new learning-based approach to tackle the challenge of joint demosaicing and denoising on image sensors.
 Wei et al.
-[35] investigate the low-light image denoising considering the sensor’s photoelectric properties.
+[31] investigated the low-light image denoising considering the sensor photoelectric properties.
 Yue et al.
-[42] achieve state-of-the-art raw image denoising by constructing a dynamic video dataset with noise-clean pairs.
+[39] achieved state-of-the-art raw image denoising by constructing a dynamic video dataset with noise-clean pairs.
 Learning-based raw image processing has demonstrated outstanding potential for high-performance reconstruction from raw sensor data.
@@
 3 New raw LDR-HDR pair Formation Pipeline
-We begin by analyzing the sensor response of the imaging system and presenting a novel data formation pipeline for raw LDR-HDR pairs based on the camera response model.
+We first analyze the sensor response of the imaging system and propose a new formation pipeline for raw HDR-paired data based on the camera response model.
 3.1 Analysis of CMOS Imaging System
@@
 where $(x,y)$ represents spatial coordinates on the sensor plane, $T$ is the integration time (exposure time), $E(x,y,\lambda)$ signifies the incident spectral irradiance, $S(x,y)$ characterizes the spatial response of the collection site, and $q(\lambda)$ is defined as the ratio (electrons/Joule) of collected electrons per incident light energy for the sensor as a function of wavelength $\lambda$.
-Given that $(x,y)$ in Eq. 1 pertains to a single photosensory cell, we assume that each parameter remains constant with respect to position. Consequently, the coordinates $(x,y)$ can be omitted [14]:
+Given that $(x,y)$ in Eq. 1 pertains to a single photosensory cell, we assume that each parameter remains constant concerning position. Consequently, the coordinates $(x,y)$ can be omitted [12]:
 […]
 where $\overline{S}$ denotes the expected value of $S(x,y)$ within a single photosensory cell, and $A$ denotes the effective photoreceptor area of the cell.
-Subsequently, the camera’s amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [47]:
+Subsequently, the camera amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [45]:
 $D=\frac{K_{\textrm{a}}Q+V_{\textrm{offset}}}{\eta}\times K_{\textrm{d}},$ (3)
@@
 We observe from Eq.
-5 that under noise-free conditions, adjusting the gain factor ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly adjust the camera’s raw response value.
-This linear characteristic allows us to achieve an equivalent result as modifying the exposure time ($T$) by simulating the Digital Gain, thereby obtaining a bracketed exposure raw image.
-Nonetheless, due to noise in real-world imaging, changing the Digital Gain will also linearly amplify the noise.
-Therefore, to mitigate noise influence, we avoid this problem as much as possible during data acquisition (in Sec.
+5 that under noise-free conditions, adjusting the gain factor ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly change the camera raw response value.
+This linear characteristic allows us to achieve an equivalent result to modifying the exposure time ($T$) by simulating the gain, thereby obtaining a set of bracketed exposure raw images.
+However, there is unavoidable noise in the actual imaging process.
+Therefore, we follow the existing denoising methods [39, 32] and try to avoid the effect of noise as much as possible during data acquisition (in Sec.
 3.2.1).
-Also, we avoid using multiplicative gain to construct long-exposure images that amplify the noise difference between long- and short-exposure images and instead use truncation to similar long-exposure images.
-In addition, we synthesize the long-short exposure images by measuring the accurate HDR sensor noise model to eliminate the long-short exposure image noise difference.
-Based on this rationale, we propose an innovative and efficient pipeline for generating LDR-HDR data pairs.
-3.2 Formation of Raw LDR-HDR Image Pair
-Relative to sRGB images, HDR reconstruction from raw images offers several distinct advantages: it retains a more significant amount of original information and has higher bit depths, which is more beneficial for improving the dynamic range of the image; the Image Signal Processing (ISP) pipeline is more straightforward and less computational demands.
-These characteristics render raw HDR reconstruction a highly promising approach for deployment in edge devices.
-Considering these considerations, we create a raw HDR dataset featuring a collection of LDR-HDR data pairs, aptly named RealRaw-HDR.
+3.2 Formation of Short- and Long-exposure Raw Pairs
+Compared to sRGB images, HDR reconstruction from raw images has the advantages of more original information, simpler ISP processing, and less computation, making it a promising paradigm to deploy in edge devices.
+To this end, we construct a new raw HDR dataset with LDR-HDR data pairs, named RealRaw-HDR.
 3.2.1 Data Acquisition
 Based on the analysis in Sec.
-3.1, we observe that changing the Digital Gain yields a comparable luminance to adjusting the exposure time when the noise level remains sufficiently low.
-Consequently, we can capture two raw images with identical exposure settings and subsequently manually simulate the Digital Gain to generate two raw images with differing exposures.
-It’s crucial to emphasize that these two raw images with the same exposure require low noise and no overexposure.
-In order to capture near-noise-free raw images, we utilize high-end cameras, special photography techniques, and post-processing acquisitions.
-Specifically, we use a top-of-the-line FUJI-FILM GFX50S II camera with a wide-aperture lens to capture high-quality raw images.
-The camera has a $51$ megapixel medium format image sensor with a pixel size of $5.3\mu m$ (The iPhone 14 Pro Max primary camera single pixel size is only $1.22\mu m$).
-We also set the camera’s ISO to 800 or below and turned on the noise reduction feature to enhance image quality.
-Subsequently, we subject the raw image to post-processing using the BM3D algorithm, attaining nearly noise-free raw images.
-Meanwhile, to simulate the relative motion between long and short exposure images within a real-world scene, we use a human subject to simulate motion between LDR images and trigger the shutter twice in a rapid time interval.
-To eliminate the risk of unintended camera shake, we securely mount the camera on a tripod and use a remote smartphone to control the shutter release.
-Following image capture, the raw images undergo black level correction and normalization before being subjected to BM3D processing [23] to effectively reduce noise, thereby yielding raw images nearly devoid of noise.
+3.1, we find that changing the sensor digital Gain, $K_{\textrm{d}}$, can achieve a similar luminance to adjusting the exposure time $K_{\textrm{a}}$ on the noise-free condition.
+Consequently, we use a top-of-the-line FUJI-FILM GFX50S II camera with a wide-aperture lens to capture high-quality raw images.
+The camera has 15.5 stops of dynamic range (15.5 bit) and a $51$ megapixel medium format image sensor with a pixel size of $5.3\mu m$ (The iPhone 15 Pro Max primary camera single pixel size is only $1.22\mu m$).
+We also set the camera ISO to 800 or below and turned on the noise reduction feature to enhance image quality.
+At this point, the captured raw image has a low noise level.
+Specifically, we capture two raw images ($I_{1}$ and $I_{2}$) with the same exposure settings using a high-end camera.
+Meanwhile, we use a human subject to simulate motion between images and trigger the shutter twice in a rapid time interval, to simulate the relative motion between short- and long-exposure images within a dual-exposure sensor.
+Further, to eliminate the risk of unintended camera shake, we mount the camera on a tripod and use a remote smartphone to control the shutter release.
+Afterward, the raw images are black level corrected, normalized, and then processed with BM3D [23] to reduce noise, which obtains nearly noise-free raw images.
+Note that there is a small relative motion between these two raw images, which is common in multi-frame HDR reconstruction.
+Our dataset will be released after the acceptance of this work.
+Figure 1: The raw LDR-HDR pair formation pipeline.
+Two clean high dynamic range raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction, and normalization.
+After manual digital gain, clip, add noise, and normalization, the long-exposure image $I_{l}$ is overexposed in the bright areas, and the short-exposure image $I_{s}$ dark area information is covered by noise.
 3.2.2 Data Processing
-Fig. 2 shows the proposed data formation pipeline. Here’s an in-depth breakdown of the data formation pipeline:
+Based on the digital camera imaging theory, we utilize two raw images ($I_{1}$ and $I_{2}$ have relative motion and are noise-free) to simulate short- and long-exposure images and construct the corresponding ground truths based on the principles of HDR synthesis.
+The Fig.
+1 shows the proposed data formation pipeline.
+Figure 2: Real samples collected by the proposed raw LDR-HDR pair formation pipeline. For display purposes, we do not apply luminance alignment processing.
 Selection of exposure time ratio and initial adjustment.
 Commencing the pipeline, we pack two clean Bayer raw images.
-We randomly select an exposure time ratio $r$ from {4, 8, 16} at this stage.
-[^1] It’s important to note that exposure ratios may not necessarily be multiples of 2.
+We select an exposure time ratio $r$ from {4, 8, 16} at this stage.
 The two normalized raw images, denoted as $I_{1}$ and $I_{2}$, multiply by the maximum pixel value (Max: $2^{12}\times r$).
@@
 Moving forward, we divide $I^{{}^{\prime}}_{1}$ by 1 and then clip the pixel values to a range of 0 to 4095 (12 bit).
-This operation is equivalent to adjusting the Digital Gain ($K{\textrm{d}}$), achieving an outcome comparable to altering the exposure time.
-Subsequently, we synthesize the noisy image by long exposure image noise model measured by Sony IMX327 HDR sensor and normalize it.
+This operation is equivalent to adjusting the sensor gain ($K{\textrm{d}}$), achieving an outcome comparable to altering the exposure time (as shown in Fig.
+2).
+Then, we add noise to create the corresponding noisy long-exposure raw images.
 This process simulates a long-exposure noisy image ($I_{l}$) reaching saturation signal level (full well capacity) and the inherent noise generated by the image sensor, which preserves dark detail while losing details in brighter areas.
@@
 Simultaneously, we divide $I^{{}^{\prime}}_{2}$ by $r$ and clip the pixel values to a range of 0 to 4095.
-(Equivalent to adjusting the Digital Gain).
-Similar to the previous step, synthesize the short-exposure noisy image by the short-exposure image noise model measured by the Sony IMX327 HDR sensor, and normalization follows suit.
+(Equivalent to adjusting the sensor gain).
+Similar to the previous step, we add noise to create the corresponding noisy short-exposure raw images.
 This procedure simulates a practical short-exposure image ($I_{s}$).
@@
 Ground truth image ($I_{gt}$).
-Additionally, we normalize $I^{{}^{\prime}}_{2}$ to obtain the ground truth image $I_{gt}$.
-The $I_{gt}$ contains more information on bright regions than $I_{l}$; $I_{gt}$ has a higher signal-to-noise ratio in dark regions compared to $I_{s}$.
+HDR aims to recover detailed information from LDR images in both brighter and dark areas.
+Therefore, for a dual-exposure HDR sensor, we aim to recover the darkest areas of the short-exposure image from the long-exposure image.
+Thus, based on this principle, we normalize $I^{{}^{\prime}}_{2}$ to obtain the ground truth image $I_{gt}$.
+The $I_{gt}$ contains more information on bright regions than $I_{l}$; $I_{gt}$ has a higher signal-to-noise ratio in dark regions than $I_{s}$.
 As a result, our data formation pipeline efficiently generates an extensive array of LDR-HDR data pairs.
@@
 Finally, after the luminance alignment [14], we obtain the noisy raw LDR images $I^{{}^{\prime}}_{l}$ and $I^{{}^{\prime}}_{s}$, and the corresponding clean raw HDR image $I^{{}^{\prime}}_{gt}$.
-The data formation pipeline facilitates the generation of rich LDR-HDR pairs, which is integral to our research and exploration.
-Figure 2: The raw LDR-HDR pair formation pipeline. Two clean high-bit raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction, normalization, and BM3D denoising.
-Figure 3: Real samples collected by the proposed raw LDR-HDR pair formation pipeline. For display purposes, we do not apply luminance alignment processing.
+Our degraded dataset follows the principle of HDR synthesis—namely, the principle of maximum signal-to-noise ratio.
+In the darkest region, we select long-exposure images; in the brightest region, we select short-exposure images.
 3.2.3 RealRaw-HDR Dataset
@@
 To the best of our knowledge, there is an absence of a raw HDR dataset explicitly tailored for these HDR sensors.
-Furthermore, it is challenging to construct large-scale, high-quality raw HDR datasets from multi-exposed frames.
-Traditional methods of collecting HDR datasets often lead to misalignment issues, yielding fused HDR images of subpar quality; previous data collection methods necessitate subjects to remain static while capturing multiple exposures, a time-consuming and labor-intensive process.
-In contrast, our proposed raw LDR-HDR pair formation pipeline is efficient and user-friendly, enabling the effortless creation of a multitude of high-quality data pairs.
-Through this pipeline, we gather $240$ pairs of $8192\times 6192$ high-resolution raw LDR-HDR image pairs.
+Our proposed data formation pipeline is efficient and user-friendly, enabling the creation of many high-quality data pairs effortlessly.
+We gather $240$ pairs of $8192\times 6192$ high-resolution raw image pairs and expand to 720 pairs.
 Fig.
@@
 1, we compare the statistics of our dataset with those of other existing HDR datasets.
-[^2] If not specified, all raw images in this paper have been processed by a fixed ISP to obtain sRGB versions for visualization purposes.
-Table 1: Comparison with existing HDR datasets.
-×$\times 3$ indicates that the same set of images can be expanded with three exposure ratios.
-Theoretically, Our dataset can be expanded $N$ times by the raw LDR-HDR pair formation pipeline, N denotes the number of covered exposure ratios.
+In this paper, all raw images have been processed with a fixed ISP, and HDR images are processed with the same tone mapping operator to obtain the sRGB version for visualization.
+Table 1: The statistics comparison between Kalanatri [14], Chen [4] and our RealRaw-HDR dataset.
 Data Quantity Size Format Exposure Ratio
-Kalanatri [16] 74 $1490\times 989$ sRGB {4, 8, 16}
-Chen [5] 144 $4096\times 2168$ raw, sRGB {4, 8, 16}
-Ours $240\times 3$ $8192\times 6192$ raw, sRGB {4, 8, 16}
-4 The Proposed RepUNet Network
+Kalanatri [14] 74 $1490\times 989$ sRGB 4 &\&& 8 &\&& 16
+Chen [4] 144 $4096\times 2168$ raw, sRGB 4 &\&& 8 &\&& 16
+Ours 720 $8192\times 6192$ raw, sRGB 4-16
+3.2.4 Effectiveness of the Data Formation Pipeline
+The proposed pipeline for generating HDR data is efficient and user-friendly, allowing easy generation of numerous high-quality data pairs.
+Although our ground truth is derived from a single image, it contains a wide range of information characteristics of HDR images.
+Firstly, the raw images $I_{1}$ and $I_{2}$ are captured by a high dynamic range camera and contain high dynamic range.
+On the other hand, there is a significant difference in the signal-to-noise ratio between HDR images and LDR images.
+The exposure-aligned long-exposure images differ from the short-exposure images only in the dark and overexposed regions, in addition to the noise difference.
+Therefore, $I_{l}$ and $I_{s}$ have all the characteristics of real-world long and short-exposure images, and $I_{gt}$ contains a wide range of informative features of real HDR images.
+4 Methodology
+4.1 Overview
 HDR reconstruction plays a vital role in various applications, such as mobile photography, high-definition displays, and virtual reality, where lightweight and efficient algorithms are highly demanded due to resource limitations.
 Previous learning-based HDR methods [21, 20, 33] often rely on large and complex models, making them impractical for real-world scenarios.
-Therefore, we propose a UNet-like lightweight base model, DualUNet, to address this challenge and employ reparameterization techniques to improve its representation capability while balancing storage and computation cost and reconstruction accuracy.
-We first present the DualUNet and then develop our RepUNet by reparameterization, followed by the design of the loss function.
-4.1 Base Model of DualUNet
-To ensure high inference speed and cross-device deployment on commodity mobile devices, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a UNet-like network as the base model, namely DualUNet.
-The overall architecture of the DualUNet is shown in Fig.
-4(a).
-We do not use the computationally demanding explicit alignment in our HDR network for lightweight design.
-We introduce an alignment-free and motion-aware short-exposure-first selection loss (in Sec.
-4.3) that enables training with unaligned pairs to compensate for the absence of alignment modules.
-We first use the pixel unshuffle operation [10] to transfer the input raw images $I^{\textquoteright}_{1}$ and $I^{\textquoteright}_{s}$ from $C\times H\times W$ to $4C\times\frac{1}{2}H\times\frac{1}{2}W$ to extract large-scale contextual information while maintaining low computational complexity.
-Notably, short-exposure images have less motion blur distortion but severe noise in dark areas.
-Conversely, long-exposure images generally contain more information with a higher signal-to-noise ratio but may suffer from significant motion blur and overexposure.
-To better leverage the complementary information from the two types of images, we introduce two distinct sub-encoders: Encoder-S and Encoder-L.
-Encoder-S extracts features from the short-exposure image, serving as foundational reference features.
+On the other hand, unlike GPU servers, existing optimized manipulations for mobile devices are quite limited, especially on computationally limited devices.
+Unsupported operators have to be processed on the CPU, which not only very low processing speed, but also introduces additional MACs.
+Therefore, we first design a neat UNet with mobile-friendly operations as the base model, then propose a re-parameterizable Topological Convolution Block to improve HDR performance.
+For lightweight design, we do not use the computationally demanding explicit alignment in our HDR network.
+To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss (in Sec.
+4.5) that enables training with unaligned pairs.
+4.2 Base Model
+To ensure high inference speed and cross-device deployment on commodity mobile devices, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a neat UNet consisting of the most basic operations as the base model.
+The overall architecture of the base model is shown in Fig.
+3(a), maximizing the use of the dual-exposure HDR sensor imaging characteristics.
+We introduce two distinct sub-encoders based on the differences in long- and short-exposure image features: Encoder-S and Encoder-L.
+Encoder-S extracts features from the short-exposure image, serving as reference features.
 In parallel, Encoder-L extracts features from the long-exposure image, offering supplementary features.
-We fuse the multi-scale features using addition or concatenation operations at each feature level to facilitate complementary features to learn motion information from the reference features.
-Each Encoder-S consists of a pixel unshuffle $\downarrow 2$ downsampling operation, a $3\times 3$ convolution layer, and a ReLU activation ([Down-Conv-ReLU]).
-At the same time, the Encoder-L has the structure of [Down-Conv-ReLU-Conv-ReLU].
-The pixel unshuffle operation can avoid information loss introduced by conventional downsampling/pooling operations.
-Finally, the fused features are fed into the decoder to generate the final HDR output.
-Refer to caption (a) DualUNet
-Refer to caption (b) Topology Convolution Block (TCB)
-Figure 4: Illustration of (a) DualUNet and (b) Topology Convolution Block (TCB).
-In the training phase, the TCB module replaces the normal convolution, which is merged back into a normal 3 × 3 convolution layer in the inference phase.
-4.2 RepUNet Network
-The representation capability of the plain DualUNet for HDR imaging is significantly limited, as shown in Tab.
-6, leading to unsatisfactory performance.
-To address this concern, we introduce a flexible re-parameterizable module called the Topological Convolution Block (TCB) and replace the normal convolution in DualUNet with the TCB module during the training phase.
-Notably, the last layer of DualUNet is still a normal convolution.
-Although the reparameterization has achieved promising results on other tasks [9, 4, 7, 46], its application in reparameterization blocks for low-level visual tasks remains scarce.
-We integrate previous reparameterization blocks and flexibly design reparameterization blocks with six operator combinations to enhance the feature representation of the model for HDR and denoising tasks.
-The TCB consists of several fundamental units: (1) a standard $3\times 3$ convolution for a solid foundation; (2) extending and squeezing convolution to enhance feature expressiveness, (3) Sobel and Laplacian operators for extracting first and second-order spatial derivatives to identify edges; (4) a jump connection to avoid gradient vanishing or exploding, and (5) a $1\times 1$ convolution to encourage information exchange between channels.
-It’s paramount to underscore that we exclusively employ the TCB with the Laplacian operator within the decoder.
+Considering the limited bandwidth, we first employ the pixel unshuffle operation [8] to transfer the input raw images $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{s}$ from $C\times H\times W$ to $4C\times\frac{H}{2}\times\frac{W}{2}$ to extract multi-scale contextual information while keeping the MAC of our model as low as possible.
+Subsequently, reference and complementary features are extracted by different numbers of normal convolutions, respectively.
+To be more specific, each layer of Encoder-S consists of a pixel unshuffle $\downarrow 2$ downsampling operation, a $3\times 3$ convolution layer, and a ReLU activation ([Down-Conv-ReLU]).
+At the same time, each layer of Encoder-L has the structure of [Down-Conv-ReLU-Conv-ReLU].
+To promote the complementary features to learn the relative motion from the reference features, we feed the reference features of each layer to the next layer by the addition of the reference features with the complementary features.
+Finally, we concatenate the reference features with the complementary features and feed them to the decoder.
+The decoder only contains 5 normal convolutions and 3 upsampling operators.
+By delicate design, the proposed base model is well-suitable for mobile scenarios with high efficiency and flexibility.
+The network design with low MAC allows for ultra-fast inference on mobile devices, and the basic operation makes cross-device deployment easier.
+(a) Base Model
+(b) Topology Convolution Block (TCB)
+Figure 3: Illustration of (a) Base Model and (b) Topology Convolution Block (TCB).
+In the training phase, the TCB employs multiple branches, which can be merged into one normal convolution layer in the inference stage.
+Figure 4: Visual comparison of state-of-the-art HDR reconstruction methods on our synthetic test dataset.
+4.3 Topological Convolution Block
+Although the plain base model is efficient, its HDR performance is less satisfactory compared to those complicated models, as shown in Tab.
+5.
+We thus employ the re-parameterization technique to enrich the representation capability of the base model.
+The reparameterization has achieved promising results on other tasks [7, 3, 5, 44].
+We design a flexible re-parameterizable module called the Topological Convolution Block (TCB)), which can more effectively extract edge and texture information for the HDR task.
+As shown in Fig.
+3(b), the TCB consists of several fundamental units: (1) A standard $3\times 3$ convolution for a solid foundation.
+The standard convolution is denoted as:
+$F_{n}=W_{n}*X+B_{n},$ (6)
+where $F_{n}$, $X$, $W_{n}$, and $B_{n}$ represent the output feature, input feature, weights, and bias of the standard convolution, respectively.
+(2) Extending and squeezing convolution to enhance feature expressiveness, the expanding and squeezing feature is extracted as:
+$F_{es}=W_{s}*(W_{e}*X+B_{e})+B_{s},$ (7)
+where $W_{e}$, $B_{e}$ and $W_{s}$, $B_{s}$ are the $1\times 1$ expanding and $3\times 3$ squeezing convolutions weights, bias, respectively.
+(3) Sobel and Laplacian operators for extracting first and second-order spatial derivatives to identify edges, i.e., using a predetermined convolution kernel to process the edge information.
+Denote by $D_{x}$ and $D_{y}$ the horizontal and vertical Sobel filters, $D_{lap}$ is the Laplacian filter.
+$D_{x}=\begin{bmatrix}+1&0&-1\\ +2&0&-2\\ +1&0&-1\end{bmatrix},$ (8)
+$D_{y}=\begin{bmatrix}+1&+2&+1\\ 0&0&0\\ -1&-2&-1\end{bmatrix},$ (9)
+$D_{lap}=\begin{bmatrix}0&+1&0\\ +1&-4&+1\\ 0&+1&0\end{bmatrix}$ (10)
+The combined edge information is extracted by:
+$F_{edge}=F_{Dx}+F_{Dy}+F_{lap},$ (11)
+where $F_{Dx}$, $F_{Dy}$, and $F_{lap}$ represent the horizontal, vertical, 2nd-order edge information, respectively.
+(4) A $1\times 1$ convolution to encourage information exchange between channels, denoted as:
+$F_{c}=W_{c}*X+B_{c},$ (12)
+where $F_{c}$, $W_{c}$, and $B_{c}$ represent the output feature, weights, and bias of the $1\times 1$ convolution, respectively.
+(5) A jump connection to avoid gradient vanishing or exploding, denoted as:
+$F_{j}=X$ (13)
+The output of the TCB in the combination of the five components:
+$F_{TCB}=F_{n}+F_{es}+F_{edge}+F_{c}+F_{j}$ (14)
+The combined feature map is then fed into a non-linear activation layer.
+PReLU is employed in our experiments.
+It is paramount to underscore that we exclusively employ the TCB with the Laplacian operator within the decoder.
 This selective approach is grounded in Laplacian operator effectiveness for noise-free images, underpinning its application to enhance feature representation in contexts devoid of noise.
+4.4 Re-parameterization for Efficient Inference
 To achieve an efficient HDR network that meets the stipulated design prerequisites of low computational complexity and streamlined hardware device deployment, we simplify the TCB reparameterization into a single $3\times 3$ convolution after training.
-Following previous works [8, 7, 46], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolution into a normal $3\times 3$ convolution.
+Following previous works [6, 5, 44], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolution into a single $3\times 3$ convolution.
 Additionally, we combine the Sobel and Laplacian operators into a special $3\times 3$ convolution with a fixed convolution kernel.
 The $1\times 1$ convolution is achieved by padding the convolution kernel with zeros.
-As a result, TCB can be transformed into a single $3\times 3$ convolution for efficient implementation during the inference stage.
+As a result, TCB can be transformed into a $3\times 3$ convolution for efficient implementation during the inference stage, as shown in Fig.
+3(b).
 By utilizing TCB, we achieve superior HDR results with improved efficiency.
-4.3 Loss Functions
-Alignment-free and motion-aware short-exposure-first selection loss.
-In fused-based HDR imaging, eliminating ghosting caused by motion inconsistencies between short and long frames is one of the most challenging issues.
-Previous work [16, 37] commonly employs optical flow, attention mechanisms, and other methods to establish pixel correspondences between long and short frames.
-The objective is to suppress ghosting by designing more elaborate fusion strategies.
-Yet, these motion estimation and alignment methods are often the most computationally intensive components and cannot be accommodated by the current level of hardware design.
-To ensure fast inference and easy hardware deployment, we abandon explicit motion alignment schemes and instead explore a more efficient alternative.
-Consequently, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate the ghost artifacts.
-Unlike other tasks involving image alignment, such as video motion estimation and stereo matching, long and short-exposure image fusion in HDR reconstruction does not necessarily require pixel-level correspondence in real-world scenes.
-The reason is that it is challenging to recover sharp object edges due to motion blur.
-In contrast, short-exposure images exhibit less motion blur distortion.
-Therefore, ghost artifacts can be suppressed by simply detecting motion regions in long-frame images through some mechanism, discarding these pixels during the fusion process, and relying solely on the corresponding regions in short-exposure images as the exclusive information source for fusion.
-Based on the same consideration, overexposed regions in the long-exposure image should likewise be discarded from the fusion process.
-Therefore, in practical engineering applications, short exposures are often used as reference images.
-Based on the above analysis, we devise the following strategies for double-exposure HDR fusion: in scenarios involving motion or overexposure within the fused region, we prefer to select the short-exposure image that contains more information; when the signal-to-noise ratio of the short frame is too low, we prefer to select the long-exposure image that contains more information.
-It is worth noting that our strategy is inclined to address motion artifacts with a higher priority than low signal-to-noise ratios.
-Consequently, when both ghost artifacts and low signal-to-noise ratios are concurrently evident, the short-exposure image is chosen to prioritize the motion artifacts.
-Figure 5: An illustrative sample of data construction for the proposed pseudo location-based motion alignment loss.
-Figure 6: Visual comparison of state-of-the-art HDR reconstruction methods on our synthetic test dataset.
-Note that after the luminance alignment, the long and short frames already have similar luminance levels, and the regions with remarkable luminance differences are mainly caused by motion and overexposure.
-Therefore, we expect the network to perceive the motion and overexposed regions by focusing on the luminance gap between the long and short exposures, thereby realizing the short-frame prioritization strategy mentioned above.
-For this purpose, we design an alignment-free and motion-aware short-exposure-first selection loss function.
-$L_{\textrm{AMSS}}=1-MSSSIM(\widetilde{I}^{out}\odot M,{I}^{gt}\odot M),$ (6)
-where $\odot$ denotes the point-wise multiplication, and $M$ is a binary mask with “1” for motion regions in long frames, and “0” otherwise.
-Given a mask $M$ indicating motion and overexposed regions, the above loss formula implements a strategy that encourages short-frame prioritization.
-To avoid explicit motion alignment, we adopt the synthetic motion to construct the mask $M$ in the training quadruplets {$I^{{}^{\prime}}_{l}$, $I^{{}^{\prime}}_{s}$, $M$, $I^{{}^{\prime}}_{gt}$.
-Specifically, as shown in Fig.
-5, for each patch of an image, we randomly select a rectangle (40 to 60-pixel range) of random length and width from the long-exposure image and then move and overlap it to a random location in the range of -30 to 30 relative to the patch.
-The patch regions before and after the movement are labeled as 1s in the mask $M$.
-By introducing the mask $M$ to guide the network to focus on moving and overexposed regions, the model effectively prioritizes the short-exposure information over the long-exposed counterpart within these regions.
-The masks are used only in training, and the inference stage inputs only short and long exposure frames.
-Reconstruction loss.
-For saturated areas, the L2 loss punishes any deviation of pixel values from the ground truth.
-This allows the model to select short-exposure information in over-exposed areas.
-To achieve the best HDR reconstruction results, we employ the MS-SSIM loss guide model, which learns short-exposure image information for global motion.
-By combining these loss functions, our model effectively produces superior results for both areas with motion and saturated regions.
-$L_{\textrm{pix}}=\|\widetilde{I}^{out}-{I}^{gt}\|_{2},$ (7)
-$L_{\textrm{ssim}}=1-MSSSIM(\widetilde{I}^{out},{I}^{gt}).$ (8)
+Figure 5: An illustrative sample of data construction for the proposed alignment-free and motion-aware short-exposure-first selection loss.
 Table 2: Performance comparison of different HDR models on our synthetic dataset.
-Params and FLOPs represent the total number of network parameters and floating-point operations.
+#Param and FLOPs represent the total number of network parameters and floating-point operations.
 The FLOPs and Run Times results are measured on an RTX 3090 device with a resolution of $4096\times 2952$ raw images.
@@
 "-" indicates the result is not available.
-Methods FLOPs Params Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16
+Methods FLOPs #Param Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16
 […]
@@
 NHDRRNet[37] 826.17G 40.26M 7.9 ms 33.0206 2.7308 33.0127 2.7277 33.0392 2.7203 33.0101 2.7443
-UNet-SID[3] 640.89G 7.76M 3.1 ms 44.3892 1.3434 44.3314 1.3535 44.4551 1.3312 44.3811 1.3456
-SGN[10] 712.66G 4.78M 3.3 ms 44.6094 1.3235 44.5074 1.3398 44.7078 1.3067 44.6131 1.3240
+UNet-SID[2] 640.89G 7.76M 3.1 ms 43.3892 1.3434 43.3314 1.3535 43.4551 1.3312 43.3811 1.3456
+SGN[8] 712.66G 4.78M 3.3 ms 43.6094 1.3235 43.5074 1.3398 43.7078 1.3067 43.6131 1.3240
 HDR-Transformer[21] 3698.28G 1.23M - 44.3895 1.3681 44.3311 1.3811 44.4140 1.3619 44.4235 1.3613
@@
 Ours 127.55G 0.82M 2.9 ms 44.8081 1.2886 44.7575 1.3000 44.8482 1.2812 44.8187 1.2842
-Figure 7: Visual comparisons with the state-of-the-art methods on HDR sensor raw dataset.
+4.5 Loss Functions
+Alignment-free and motion-aware short-exposure-first selection loss.
+In fused-based HDR methods, eliminating ghosting caused by motion inconsistencies between short- and long-exposure pairs is one of the most challenging issues.
+Previous work [14, 34] commonly employs optical flow, attention mechanisms, and other methods to establish pixel correspondences between short- and long-exposure images.
+The objective is to suppress ghosting by designing more elaborate fusion strategies.
+However, these motion estimation and alignment methods are often the most computationally intensive components and cannot be accommodated by the current level of hardware design.
+On the other hand, unlike other image alignment tasks, such as video motion estimation and stereo matching, short- and long-exposure image fusion in HDR reconstruction does not necessarily require pixel-level correspondence.
+The reason is that it is challenging to recover sharp object edges due to motion blur.
+In contrast, short-exposure images exhibit less motion blur distortion.
+Therefore, ghost artifacts can be suppressed by simply detecting motion regions in long-frame images through some mechanism, discarding these pixels during the fusion process, and relying solely on the corresponding regions in short-exposure images as the exclusive information source for fusion.
+Based on the same consideration, overexposed regions in the long-exposure image should likewise be discarded in the fusion process.
+Short-exposure images are often used as reference images in engineering applications.
+Based on the above analysis, we devise the strategies for dual-exposure HDR fusion: Firstly, in scenarios involving motion or overexposure within the fused region, we prefer to select short-exposure image that contains more information; Secondly, when the SNR of the short frame is too low, we prefer to select the long-exposure image that contains more information; Thirdly, our strategy is inclined to address ghost artifacts with a higher priority than lower SNR when ghost artifacts and lower SNR concurrently exist.
+For the designed fusion strategy, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate the ghost artifacts.
+We first construct a mask $M$ in the training pairs {$I^{{}^{\prime}}_{l}$, $I^{{}^{\prime}}_{s}$}.
+Specifically, as shown in Fig.
+5, for each patch of an image, we randomly select a rectangle of random length and width from the long-exposure image and then move and overlap it to a random location in the range of -30 to 30 relative to the patch.
+The patch regions before and after the movement are labeled as 1s in the mask $M$.
+By introducing the mask $M$ to guide the network to focus on moving and overexposed regions, the model effectively prioritizes the short-exposure information over the long-exposed counterpart within these regions.
+The masks are used only in training, and the inference stage inputs only short and long exposure frames.
+The alignment-free and motion-aware short-exposure-first selection loss is denoted as:
+$L_{\textrm{AMSS}}=1-\text{MS-SSIM}(\widetilde{I}^{out}\odot M,{I}^{gt}\odot M),$ (15)
+where $\odot$ denotes the point-wise multiplication, and $M$ is a binary mask with “1” for motion regions in long frames, and “0” otherwise, MS-SSIM denotes multi-scale structural similarity function.
+Given a mask $M$ indicating motion and overexposed regions, the above loss formula implements a strategy that encourages short-frame prioritization.
+Reconstruction loss. For saturated areas, the L2 loss punishes any deviation of pixel values from the ground truth. This allows the model to select short-exposure information in over-exposed areas.
+$L_{\textrm{pix}}=\|\widetilde{I}^{out}-{I}^{gt}\|_{2},$ (16)
+To achieve the best HDR reconstruction results, we employ the multi-scale structural similarity loss function loss guide model, which learns short-exposure image information for global motion.
+By combining these loss functions, our model effectively produces superior results for both areas with motion and saturated regions.
+$L_{\textrm{ssim}}=1-\text{MS-SSIM}(\widetilde{I}^{out},{I}^{gt})$ (17)
 Bayer loss.
@@
 where $\alpha$, $\beta$, $\gamma$, and $\eta$ are the corresponding weight coefficients.
-5 Experiments
-5.1 Experimental Setup
-Datasets and Metrics.
-We utilize the proposed RealRaw-HDR dataset for training.
-For testing, we first evaluate our method on the dataset, which is synthesized using the data formation pipeline.
-This dataset was not utilized during training.
-This test set contains 30 sets of test samples containing different exposure ratios (i.e., 4, 8, and 16) with a resolution of $4096\times 2176$.
-To validate the generalization ability of our method on real data, we utilize a FUJI-FILM GFX50S II camera to capture seven sets of real-world bracketed exposure raw images and the corresponding static images for generating the ground truth.
-Furthermore, to validate the generalization ability of our method on HDR sensors, we utilize Chen’s test dataset [5] for cross-validation, which has short-long raw images captured by a Sony IMX267 image sensor.
-We perform a quantitative evaluation using the PSNR, SSIM, and CIE L*a*b* space  [^3] CIE L*a*b* is a color space specified by the International Commission on Illumination.
-It describes all the colors visible to the human eye and was created to serve as a device-independent model to be used as a reference.
-[45, 13, 20] (also known as $\Delta E$).
-$\Delta E$ can effectively assess chromaticity, contrast, and color accuracy variations within HDR images rather than exclusively concentrating on luminance differences (HDR-VDP-2).
-It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE Lab color space.
-$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (11)
-where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE Lab version of the predicted HDR image and ground truth, respectively.
+Figure 6: Visual comparisons with the state-of-the-art methods on HDR sensor raw dataset.
 Table 3: Performance comparison of different HDR models on the actual HDR sensor raw dataset [4]. The best and second-best performances are in bold and underlined, respectively.
@@
 Ours 40.5238 0.9747 1.9568 40.4061 0.9733 1.9743 41.4010 0.9788 1.7974 39.7642 0.9721 2.0988
-Figure 8: Visual comparison of state-of-the-art HDR reconstruction methods on RealRaw-HDR dataset with realistic exposure ratios.
+5 Experiments
+5.1 Experimental Setup
+Datasets and metrics.
+We utilize the proposed RealRaw-HDR dataset for training.
+We first evaluate our method in the synthesized dataset.
+This test set contains 30 samples containing different exposure ratios (i.e., 4, 8, and 16) with a resolution of $4096\times 2176$.
+To validate the validity of our method on real data, we utilize a FUJI-FILM GFX50S II camera to capture seven sets of real-world bracketed exposure raw images and the corresponding static images for generating the ground truth.
+Furthermore, we also utilize the Chen [4] test dataset for cross-validation, which has short- and long-exposure raw pairs captured by a Sony IMX267 image sensor.
+We perform a quantitative evaluation using the PSNR, SSIM, and CIE L*a*b* space  [^1] CIE L*a*b* is a color space specified by the International Commission on Illumination.
+It describes all the colors visible to the human eye and was created to serve as a device-independent model for reference.
+[43, 11, 19] (also known as $\Delta E$).
+$\Delta E$ can effectively assess chromaticity, contrast, and color accuracy variations within HDR images rather than exclusively concentrating on luminance differences (HDR-VDP-2).
+It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE L*a*b* color space.
+$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (20)
+where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE L*a*b* version of the predicted HDR image and ground truth, respectively.
 Implementation details.
@@
 Our model is implemented in PyTorch and trained with an NVIDIA RTX 3090 GPU.
-5.2 Comparison with the State-of-the-art Methods
+5.2 Comparison with the Other Methods
 We choose several representative low-level vision methods for comparisons, including four HDR methods based on sRGB images (AHDRNet [34], DeepHDR [33], NHDRRNet [37], HDR-Transformer[21]), as well as two methods for denoising raw images (SGN [8] and UNet-SID [2]).
@@
 Additionally, for AHDRNet, DeepHDR, HDR-Transformer, and NHDRRNet, we modify the network inputs to accommodate dual-exposure raw images.
-Similarly, for SGN and UNet-SID, we concatenate the long- and short-exposure images as inputs.
-Table 4: Quantitative comparisons of different loss functions. AMSS-Loss represents the alignment-free and motion-aware short-exposure-first selection loss.
-ID Method Bayer-Loss AMSS-Loss All-Exposure Ratio=4 Ratio=8 Ratio=16
-[…]
-1 RepUNet 39.6403 2.1479 39.5214 2.1577 40.5216 1.9603 38.8780 2.3256
-2 RepUNet 40.0251 2.1162 39.9900 2.1143 40.8125 1.9700 39.2729 2.2642
-3 RepUNet 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
+Similarly, for SGN and UNet-SID, we concatenate the long- and short-exposure raw pairs as inputs.
 Evaluation on synthetic dataset.
@@
 In comparison, DeepHDR [33], NHDRRNet [37], and SGN [8] exhibit numerous artifacts in the palm motion region.
-However, our proposed RepUNet can reconstruct HDR images without ghosting (see rows 2 and 4).
-Evaluation on HDR sensor raw dataset.
-To validate the generalization ability of our method on the real-world HDR sensor, we utilize Chen’s test dataset [5] for cross-validation, which has raw images captured by a Sony IMX267 image sensor.
-Compared with previous methods, our method achieves state-of-the-art performance in terms of both visual quality and quantitative metrics.
+However, our proposed RepUNet can reconstruct HDR images without ghosting (see rows 2 in Fig.
+4).
+Figure 7: Visual comparison of state-of-the-art HDR reconstruction methods on FUJI raw dataset.
+Table 4: Quantitative comparisons of different loss functions. AMSS-Loss represents the alignment-free and motion-aware short-exposure-first selection loss.
+ID Method Bayer-Loss AMSS-Loss All-Exposure Ratio=4 Ratio=8 Ratio=16
+[…]
+1 RepUNet 39.6403 2.1479 39.5214 2.1577 40.5216 1.9603 38.8780 2.3256
+2 RepUNet 40.0251 2.1162 39.9900 2.1143 40.8125 1.9700 39.2729 2.2642
+3 RepUNet 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
+Table 5: Reparameterization ablation results. The FLOPs and run times are measured on the raw image with a 4K resolution.
+Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
+[…]
+Base Model 93.26G 0.82M 3.0 ms 39.7941 2.1202 39.8026 2.1036 40.5092 1.9790 39.0705 2.2780
+RepUNet 93.26G 0.82M 2.9 ms 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
+Evaluation on HDR sensor dataset.
+To validate the validity of our method on the real-world HDR sensor dataset, we utilize the Chen [4] test dataset for cross-validation, which has raw images captured by a Sony IMX267 image sensor.
+Compared with previous methods, our method achieves state-of-the-art performance in visual quality and quantitative metrics.
 The visual results from tests on the HDR sensor raw dataset (as shown in Fig.
-7) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from severe color casts.
+6) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from severe color casts.
 Furthermore, results in Tab.
 3 reveal that compared to AHDRNet [34], our method yields an improvement of more than 0.35 dB and 0.05 gain in PSNR and $\Delta E$, respectively, for scenes with an exposure ratio of 8.
-On average, our method attains gains exceeding 0.1 dB and 0.15 in these metrics.
-Evaluation on FUJI raw dataset.
-We then evaluate our method on the FUJI raw datasets, which are real-world bracketed exposure raw images captured by the FUJI-FILM GFX50S II camera.
-Fig.
-8 compares results from two high dynamic range scenes, where our method achieves significantly better visualization.
-Our method can recover both fine details in overexposed regions and rich colors in underexposed areas without introducing artifacts (see rows 1 and 2).
-Compared to AHDRNet, our method can effectively remove noise and preserve the structure of dark regions.
-Notably, the alignment module in DeepHDR, AHDRNet, HDR-Transformer, and NHDRRNet requires many line buffers, making it challenging to deploy on resource-limited edge devices.
-Notably, HDR-Transformer fails to perform inference even on RTX 3090 devices.
-In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images, as evident in the visual results of Fig.
-6, Fig.
-7, and Fig.
-8.
+On average, our method attains gains exceeding 0.1 dB and 0.15 in PSNR and $\Delta E$.
 Table 6: Performance comparison of different HDR models on RealRaw-HDR dataset with realistic exposure ratios.
@@
 Ours 475.83G 42.5364 0.9760 1.4101
-Table 6: Reparameterization ablation results. The FLOPs and Run Times are measured on the raw image of 4K resolution.
-Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
-[…]
-DualUNet 93.26G 0.82M 3.0 ms 39.7941 2.1202 39.8026 2.1036 40.5092 1.9790 39.0705 2.2780
-[…]
-RepUNet 93.26G 0.82M 2.9 ms 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
-Despite the fact that all models are trained on the RealRaw-HDR dataset, which is synthesized using the data formation pipeline, they consistently excel on both the synthetic test set and the real-world dataset.
-Particularly noteworthy is the remarkable performance achieved on the test set comprised of raw images captured by the HDR sensor [5].
-These results serve as solid evidence of the generalizability of our proposed RealRaw-HDR dataset and the HDR data formation pipeline.
-It underscores the robustness inherent in our approach and its capacity to tackle real-world challenges.
-This not only validates the efficacy of our methodology but also highlights its considerable potential for diverse real-world applications.
+Evaluation on FUJI raw dataset.
+We then evaluate our method on the FUJI raw datasets, which are real-world bracketed exposure raw images captured by the FUJI-FILM GFX50S II camera.
+Fig.
+7 compares results from two high dynamic range scenes, where our method achieves significantly better visualization.
+Our method can recover both fine details in overexposed regions and rich colors in underexposed areas without introducing artifacts (see rows 1 and 2).
+Compared to AHDRNet, our method can effectively remove noise and preserve the structure of dark regions.
+Notably, the alignment module in DeepHDR, AHDRNet, HDR-Transformer, and NHDRRNet requires many line buffers, making it challenging to deploy on resource-limited edge devices.
+HDR-Transformer fails to perform inference even on RTX 3090 devices.
+In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images, as evident in the visual results of Fig.
+4, Fig.
+6, and Fig.
+7.
+Although all models are trained on the RealRaw-HDR dataset, which is synthesized using the data formation pipeline, they consistently excel on both the synthetic test dataset and the real-world dataset.
+Particularly noteworthy is the remarkable performance achieved on the test dataset comprised of raw images captured by the HDR sensor [4].
+These results are solid evidence of the generalizability of our proposed RealRaw-HDR dataset and the HDR data formation pipeline.
 5.3 Ablation Study
@@
 To demonstrate such generalization, we transform the collected RealRaw-HDR dataset with a fixed ISP pipeline into the sRGB color space, named the Raw2RGB-HDR dataset.
-For comparison, we train the sRGB HDR method AHDRNet [37] on our Raw2RGB-HDR dataset (240 pairs) and Kalantari’s dataset [16] (taking the first two exposures as input, 74 pairs of images), respectively.
-The test dataset is from Kalantari’s dataset.
+For comparison, we train the sRGB HDR method AHDRNet [34] on our Raw2RGB-HDR dataset and Kalantari dataset [14] (taking the first two exposures as input, 74 pairs of images), respectively.
+The test dataset is from the Kalantari dataset.
 Results in Tab.
-7 show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on Kalantari’s dataset by 2.86 dB in PSNR.
+7 show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on the Kalantari dataset by 2.86 dB in PSNR.
 The performance gains benefit from an efficient and user-friendly data acquisition pipeline that generates more trainable data pairs.
 The results demonstrate that our data pipeline is also effective in generating paired LDR-HDR data in sRGB space.
-Table 7: We train the sRGB HDR method AHDRNet on our Raw2RGB-HDR dataset (240 pairs) and Kalantari’s dataset, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR dataset.
+Table 7: We train the sRGB HDR method AHDRNet on our Raw2RGB-HDR dataset and Kalantari dataset, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR dataset.
 Method Dataset PSNR PSNR-$\mu$
-AHDRNet Kalantari 35.4581 38.1618
+AHDRNet[34] Kalantari 35.4581 38.1618
 Raw2RGB-HDR 38.3183 39.8896
+Figure 8: Visual results of RepUNet and its baseline variants. Combining these loss functions allows our model to produce top-notch results for motion and saturated areas effectively.
 Loss functions.
@@
 Tab.
-4 and 9 show that adding the AMSS loss and Bayer loss steadily improves visual quality and quantitative results.
+4 and Fig.
+8 show that adding the AMSS and Bayer loss steadily improves visual quality and quantitative results.
 RepUNet with joint loss achieves the best results, outperforming the baseline by 0.5 dB in PSNR and by 0.16 in $\Delta E$ on average.
@@
 Meanwhile, our proposed Bayer loss can alleviate the color cast (see columns 2 and 3).
-Figure 9: Visual results of RepUNet and its baseline variants. By combining these loss functions, our model can effectively produce top-notch results for both motion and saturated areas.
 […]
-6 Discussion and Future Work
-Although our method can produce high-quality HDR images in real scenes, it has several limitations that may inspire future work.
-Our data synthesis pipeline employs noise simulation to synthesize long and short-exposure images.
-However, we have found that if the noise distribution of the images we test is significantly different from the noise model we have measured, the results of the reconstruction are often unsatisfactory.
-Nevertheless, for practical engineering applications, it is important to accurately measure each type of image sensor noise model to achieve efficient raw denoising.
-Furthermore, our work does not cover scenarios involving flicker.
-This is because the flicker problem generally arises in short-exposure images, and it is difficult to determine which region is better when flicker and motion occur simultaneously.
-Unfortunately, there is no clear-cut solution that can effectively address scenes where both flicker and motion are present.
-To overcome this challenge, it may be necessary to create a flicker model to generate a substantial amount of training data.
 6 Conclusion
-We propose a novel data formation pipeline for raw LDR-HDR pairs based on HDR sensor imaging systems and construct a new high-quality raw dataset and raw LDR-HDR pairs dataset.
-This dataset can efficiently support many kinds of dual-exposure HDR sensor development, which provides vital data support for realizing the new generation of AI ISPs.
-Meanwhile, we develop a lightweight and efficient RepUNet network that uses structural reparameterization for HDR image reconstruction.
-Our evaluation has shown that this model can be easily deployed on mobile devices.
-We propose a plug-and-play alignment-free and motion-aware short-exposure-first selection loss and Bayer loss to address the issue of ghosting artifacts and color cast.
-Extensive experiments demonstrate that our approach achieves state-of-the-art performance in both visual quality and quantitative metrics.
+In the paper, we proposed a Topological Convolution Block (TCB) for efficient and light-weight HDR design for mobile devices.
+Based on the proposed TCB, we further designed RepUNet, aiming at balancing hardware efficiency and PSNR/SSIM indexes.
+Furthermore, We propose a novel computational photography based pipeline for raw HDR image formation and construct a real-world raw HDR dataset – RealRaw-HDR.
+Meanwhile, we designed plug-and-play alignment-free and motion-aware short-exposure-first selection loss efficient mitigate ghost artifacts.
+Our empirical evaluation validates the effectiveness of the proposed LDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.

### v6

@@
 Abstract
-The widespread usage of high-definition screens on edge devices stimulates a strong demand for efficient high dynamic range (HDR) algorithms.
-However, many existing HDR methods either deliver unsatisfactory results or consume too much computational and memory resources, hindering their application to high-resolution images (usually with more than 12 megapixels) in practice.
-In addition, existing HDR dataset collection methods often are labor-intensive.
-In this work, in a new aspect, we discover an excellent opportunity for HDR reconstructing directly from raw images and investigating novel neural network structures that benefit the deployment of mobile devices.
-Our key insights are threefold: (1) we develop a lightweight-efficient HDR model, RepUNet, using the structural re-parameterization technique to achieve fast and robust HDR; (2) we design a new computational raw HDR data formation pipeline and construct a real-world raw HDR dataset, RealRaw-HDR; (3) we propose a plug-and-play motion alignment loss to mitigate motion ghosting under limited bandwidth conditions.
-Our model contains less than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU.
-While being highly efficient, our model also outperforms the state-of-the-art HDR methods in terms of PSNR, SSIM, and a color difference metric.
-1 Introduction
-Most resource-constrained cameras exhibit low dynamic ranges (LDR), rendering them unable to capture the full range of brightness and color information in real-world scenes.
-Conversely, high dynamic range (HDR) imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in LDR images.
+The growing prevalence of high-resolution displays on edge devices has created a pressing need for efficient high dynamic range (HDR) imaging algorithms.
+However, most existing HDR methods either struggle to deliver satisfactory visual quality or incur high computational and memory costs, limiting their applicability to high-resolution inputs (typically exceeding 12 megapixels).
+Furthermore, current HDR dataset collection approaches are often labor-intensive and inefficient.
+In this work, we explore a novel and practical solution for HDR reconstruction directly from raw sensor data, aiming to enhance both performance and deployability on mobile platforms.
+Our key insights are threefold: (1) we propose RepUNet, a lightweight and efficient HDR network leveraging structural re-parameterization for fast and robust inference; (2) we design a new computational raw HDR data formation pipeline and construct a new raw HDR dataset, RealRaw-HDR; (3) we design a plug-and-play motion alignment loss to suppress ghosting artifacts under constrained bandwidth conditions effectively.
+Our model contains fewer than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU.
+While being highly efficient, our model also achieves comparable performance to state-of-the-art HDR methods in terms of PSNR, SSIM, and a color difference metric.
+I Introduction
+Most resource-constrained cameras exhibit standard dynamic range (SDR), rendering them unable to capture the full range of brightness and color information in real-world scenes.
+Conversely, high dynamic range (HDR) imaging seeks to encompass a significantly broader range of luminance values, compensating for color distortions and the subtle detail loss observed in SDR images.
 Despite the availability of dedicated hardware for directly acquiring HDR images, such equipment is typically expensive, thus limiting its practicality for most users.
 As a result, there has been an increasing focus on fusion-based HDR imaging methods.
-Recent methods [35, 20, 4, 25, 34] based on convolution neural networks (CNNs) [17, 22] have made impressive progress in HDR reconstruction performance, thanks to their scalability and flexibility from constructing elementary building blocks like convolutional layers.
-However, superior performance is usually obtained at a cost of heavy computational burden [20, 34, 21].
+Recent methods [47, 24, 5, 31, 46] based on convolution neural networks (CNNs) [22, 51, 40, 16] have made impressive progress in HDR reconstruction performance, thanks to their scalability and flexibility from constructing elementary building blocks like convolutional layers.
+However, superior performance is usually obtained at a cost of a heavy computational burden [24, 46, 25].
 Although this can be alleviated by elaborate network structures or dedicated computing engines (e.g., GPU and NPU), the hardware cost and power consumption still limit the deployment of existing deep HDR reconstruction networks.
@@
 They first make the subject static and take three sets of images with different exposures, and then make the subject move twice to take dynamic images with different exposures.
-However, this process is labor intensive and difficult to acquire on a large scale.
-2) ISP Processing Speed: When obtaining raw LDR images with different exposures, the ISP pipeline must be performed separately on each exposure.
+However, this process is labor-intensive and difficult to acquire on a large scale.
+2) ISP Processing Speed: When obtaining raw SDR images with different exposures, the ISP pipeline [45, 9] must be performed separately on each exposure.
 This incurs additional memory and computational overhead and leads to lower frame rates for HDR image output.
@@
 The limitations of current HDR reconstruction methods highlight the need for further research and development in HDR imaging.
-In this paper, we propose an efficient scheme for HDR image reconstruction in the raw image domain.
+In this paper, we propose an efficient scheme for HDR reconstruction in the raw image domain.
 By analyzing the HDR image sensor system, we design a lightweight and efficient model for raw HDR reconstruction named RepUNet.
-RepUNet adopts reparameterization techniques and does not contain explicit computationally expensive alignment modules, such as optical flow [14], deformation convolution [4], or attention [20, 37], which are commonly used in existing deep learning-based HDR reconstruction methods [14, 20, 33, 24].
+RepUNet adopts reparameterization techniques and does not contain explicit computationally expensive alignment modules, such as optical flow [18], deformation convolution [5], or attention [24, 49, 39], which are commonly used in existing deep learning-based HDR reconstruction methods [18, 24, 44, 30].
 To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss, which encourages the network to focus on local motion patterns and alleviate misalignment between short- and long-exposure images.
@@
 This automatic control operable system satisfies real-world scenes’ dynamic range requirements, making it a practical tool for generating high-quality HDR images.
-The resulting RealRaw-HDR dataset includes many LDR-HDR pairs for training and evaluation.
+The resulting RealRaw-HDR dataset includes many SDR-HDR pairs for training and evaluation.
 By incorporating the unique characteristics of raw images into our approach, we can achieve superior HDR reconstruction results with increased efficiency and accuracy.
@@
   • $\bullet$
-    We investigate the structure re-parameterizable technique for the HDR task and propose a lightweight model, RepUNet, with Topological Convolution Block (TCB).
+    We investigate the structure reparameterizable technique for the HDR task and propose a lightweight model, RepUNet, with Topological Convolution Block (TCB).
 TCB can be used to improve the HDR performance of any HDR model without introducing any extra burden for inference.
@@
 Our contributions represent a significant step forward in raw HDR image reconstruction research, providing an effective and efficient solution for producing high-quality HDR images.
-Our model contains less than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU.
+Our model contains fewer than 830K parameters and takes less than 3 ms to process an image of 4K resolution using one RTX 3090 GPU.
 While highly efficient, our model also outperforms the state-of-the-art HDR methods by a large margin in terms of PSNR, SSIM, and a color difference metric.
-2 Related Work
-2.1 HDR Imaging
+II Related Work
+II-A HDR Imaging
 Recently, benefiting from the fast development of deep learning techniques, training deep neural networks for effective HDR reconstruction has become increasingly popular.
-Many methods apply deep neural networks [16, 24, 29, 13] to learn the production of high-quality HDR images from a set of bracketed exposure LDR images.
+Many methods apply deep neural networks [20, 30, 35, 17] to learn the production of high-quality HDR images from a set of bracketed exposure SDR images.
 Kalantari et al.
-[14] proposed a CNN-based HDR approach that employs optical flow to align LDR sRGB images before network inference.
+[18] proposed a CNN-based HDR approach that employs optical flow to align SDR sRGB images before network inference.
 Wu et al.
@@
 Yan et al.
-[34] introduced spatial attention to achieve LDR image alignment.
+[46] introduced spatial attention to achieve SDR image alignment.
 Liu et al.
 [24] presented an attention-guided deformable convolutional network for multi-frame HDR imaging.
-Prabhakar et al.
-[27] introduced an efficient method for generating HDR images using a bilateral guided up-sampler and exploring zero-learning for HDR reconstruction.
 Niu et al.
@@
 [25] proposed a Transformer-based [32] HDR imaging method.
-These deep learning-based approaches [34, 20, 21] consistently elevated state-of-the-art performance.
+SCTNet [37] proposed an alignment-free architecture employing semantic consistency transformers.
+These deep learning-based approaches [24, 25] have consistently pushed the boundaries of state-of-the-art performance.
+To address the need for deployment on mobile devices, Prabhakar et al.
+[33] introduced an efficient method for generating HDR images using a bilateral guided up-sampler and exploring zero-learning for HDR reconstruction.
+CEN-HDR introduced a lightweight architecture designed for real-time mobile applications.
 However, these methods reconstruct HDR based on sRGB images at the end of the ISP pipeline and only train on one dataset.
@@
 To simplify the ISP system, another class of HDR reconstruction methods is based on raw image input.
-Google HDR+ produced the raw HDR image by aligning and merging a burst of raw frames with the same low exposure.
+Google HDR+ [11] produced the raw HDR image by aligning and merging a burst of raw frames with the same low exposure.
 Nevertheless, this approach requires a complex ISP system design and takes up a lot of DDR memory.
@@
 Therefore, none of the existing methods can meet the requirements of real scenarios.
-2.2 Low-level Raw Image Processing
+II-B Low-level Raw Image Processing
 Due to the merits of raw data, raw-based image processing [21, 53] has made significant progress in recent years.
-The work in [40] first performs the demoireing task in the raw domain and then utilizes a pretrained ISP module to transform the result into the sRGB domain.
+The work in [54] first performs the demoiréing task in the raw domain and then utilizes a pre-trained ISP module to transform the result into the sRGB domain.
 Yang et al.
@@
 Wei et al.
-[31] investigated the low-light image denoising considering the sensor photoelectric properties.
+[43] investigated the low-light image denoising considering the image sensor photoelectric properties.
 Yue et al.
 [52] achieved state-of-the-art raw image denoising by constructing a dynamic video dataset with noise-clean pairs.
+Zou et al.
+[60] proposed a model tailor-made for Raw images, harnessing the unique features of Raw data to facilitate the Raw-to-HDR mapping.
 Learning-based raw image processing has demonstrated outstanding potential for high-performance reconstruction from raw sensor data.
 However, acquiring paired data in the raw domain is difficult and expensive.
-Our work proposes a new large-scale, high-quality raw dataset and provides a pipeline to acquire raw LDR-HDR pairs based on the imaging system.
-3 New raw LDR-HDR pair Formation Pipeline
+Our work proposes a new large-scale, high-quality raw dataset and provides a data synthesis pipeline to acquire raw SDR-HDR pairs based on the HDR sensor imaging system.
+III New Dataset Formation Pipeline
 We first analyze the sensor response of the imaging system and propose a new formation pipeline for raw HDR-paired data based on the camera response model.
-3.1 Analysis of CMOS Imaging System
+III-A Analysis of CMOS Imaging System
 The essence of a CMOS image sensor is photo-electric signal conversion.
 For a single pixel, the number of electrons $Q$ released during the light-electric signal conversion can be ideally expressed as [13]:
-[…]
+$Q=T\int_{\lambda}\int_{x}\int_{y}E(x,y,\lambda)S(x,y)q(\lambda)dxdyd\lambda,$ (1)
 where $(x,y)$ represents spatial coordinates on the sensor plane, $T$ is the integration time (exposure time), $E(x,y,\lambda)$ signifies the incident spectral irradiance, $S(x,y)$ characterizes the spatial response of the collection site, and $q(\lambda)$ is defined as the ratio (electrons/Joule) of collected electrons per incident light energy for the sensor as a function of wavelength $\lambda$.
 Given that $(x,y)$ in Eq. 1 pertains to a single photosensory cell, we assume that each parameter remains constant concerning position. Consequently, the coordinates $(x,y)$ can be omitted [15]:
-[…]
+$Q=T\overline{S}A\int_{\lambda}E(\lambda)q(\lambda)d\lambda,$ (2)
 where $\overline{S}$ denotes the expected value of $S(x,y)$ within a single photosensory cell, and $A$ denotes the effective photoreceptor area of the cell.
-Subsequently, the camera amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [45]:
+Subsequently, the camera amplifier circuit amplifies the electrical signal, yielding the raw camera response value through analog-to-digital conversion [38, 58]:
 $D=\frac{K_{\textrm{a}}Q+V_{\textrm{offset}}}{\eta}\times K_{\textrm{d}},$ (3)
@@
 $\eta$ corresponds to the quantization step associated with the bit depth.
-Combining Eq. 2 and Eq. 3, the ideal model for the optical-to-digital conversion is modeled as:
-[…]
+Combining Eq.
+2 and Eq.
+3, the ideal model for the optical-to-digital conversion is modeled as:
+$D=\frac{K_{\textrm{a}}T\overline{S}A\int_{\lambda}E(\lambda)q(\lambda)d\lambda+V_{\textrm{offset}}}{\eta}\times K_{\textrm{d}},$ (4)
 where $D$ signifies the pixel value in the raw image, $V_{\textrm{offset}}/{\eta}$ accommodates artificially introduced bias voltage to prevent output signals below 0.
 The raw response value of the bias voltage (i.e., black level) can be directly read out.
-When the dark current is 00, or we subtract the raw response value of the bias voltage, the pixel value in the raw image can be expressed as:
-[…]
+When the dark current is 0, or we subtract the raw response value of the bias voltage, the pixel value can be expressed as:
+$D=\frac{K_{\textrm{a}}T\overline{S}A\int_{\lambda}E(\lambda)q(\lambda)d\lambda}{\eta}\times K_{\textrm{d}}.$ (5)
 We observe from Eq.
-5 that under noise-free conditions, adjusting the gain factor ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly change the camera raw response value.
+5 that under noise-free conditions, adjusting the gain ($K_{\textrm{a}},K_{\textrm{d}}$) can linearly change the camera raw response value.
 This linear characteristic allows us to achieve an equivalent result to modifying the exposure time ($T$) by simulating the gain, thereby obtaining a set of bracketed exposure raw images.
 However, there is unavoidable noise in the actual imaging process.
-Therefore, we follow the existing denoising methods [39, 32] and try to avoid the effect of noise as much as possible during data acquisition (in Sec.
-3.2.1).
-3.2 Formation of Short- and Long-exposure Raw Pairs
+Therefore, we follow the existing denoising methods [52, 42] and try to avoid the effect of noise as much as possible during data acquisition.
+III-B Formation of Short- and Long-exposure Raw Pairs
 Compared to sRGB images, HDR reconstruction from raw images has the advantages of more original information, simpler ISP processing, and less computation, making it a promising paradigm to deploy in edge devices.
-To this end, we construct a new raw HDR dataset with LDR-HDR data pairs, named RealRaw-HDR.
-3.2.1 Data Acquisition
+To this end, we construct a new raw HDR dataset with SDR-HDR data pairs, named RealRaw-HDR.
+III-B1 Data Acquisition
 Based on the analysis in Sec.
-3.1, we find that changing the sensor digital Gain, $K_{\textrm{d}}$, can achieve a similar luminance to adjusting the exposure time $K_{\textrm{a}}$ on the noise-free condition.
+III-A, we find that changing the sensor digital Gain, $K_{\textrm{d}}$, can achieve a similar luminance to adjusting the exposure time $K_{\textrm{a}}$ on the noise-free condition.
 Consequently, we use a top-of-the-line FUJI-FILM GFX50S II camera with a wide-aperture lens to capture high-quality raw images.
@@
 Our dataset will be released after the acceptance of this work.
-Figure 1: The raw LDR-HDR pair formation pipeline.
-Two clean high dynamic range raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction, and normalization.
-After manual digital gain, clip, add noise, and normalization, the long-exposure image $I_{l}$ is overexposed in the bright areas, and the short-exposure image $I_{s}$ dark area information is covered by noise.
-3.2.2 Data Processing
+Figure 1: The raw SDR-HDR pair formation pipeline.
+Two clean HDR raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction and normalization.
+After manual digital gain, clip, add noise, and normalization, the long-exposure image $I_{l}$ is overexposed in the bright areas, and the short-exposure image $I_{s}$ has dark area information covered by noise.
+III-B2 Data Processing
 Based on the digital camera imaging theory, we utilize two raw images ($I_{1}$ and $I_{2}$ have relative motion and are noise-free) to simulate short- and long-exposure images and construct the corresponding ground truths based on the principles of HDR synthesis.
@@
 1 shows the proposed data formation pipeline.
-Figure 2: Real samples collected by the proposed raw LDR-HDR pair formation pipeline. For display purposes, we do not apply luminance alignment processing.
+Figure 2: Real samples collected by the proposed raw SDR-HDR pair formation pipeline. For display purposes, we do not apply luminance alignment processing.
 Selection of exposure time ratio and initial adjustment.
@@
 We select an exposure time ratio $r$ from {4, 8, 16} at this stage.
-The two normalized raw images, denoted as $I_{1}$ and $I_{2}$, multiply by the maximum pixel value (Max: $2^{12}\times r$).
+The two normalized raw images, denoted as $I_{1}$ and $I_{2}$, are multiplied by the maximum pixel value (Max: $2^{12}\times r$).
 This operation yields $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{2}$ correspondingly.
@@
 Ground truth image ($I_{gt}$).
-HDR aims to recover detailed information from LDR images in both brighter and dark areas.
+HDR aims to recover detailed information from SDR images in both brighter and darker areas.
 Therefore, for a dual-exposure HDR sensor, we aim to recover the darkest areas of the short-exposure image from the long-exposure image.
@@
 The $I_{gt}$ contains more information on bright regions than $I_{l}$; $I_{gt}$ has a higher signal-to-noise ratio in dark regions than $I_{s}$.
-As a result, our data formation pipeline efficiently generates an extensive array of LDR-HDR data pairs.
+As a result, our data formation pipeline efficiently generates an extensive array of SDR-HDR data pairs.
 Luminance alignment.
-Finally, after the luminance alignment [14], we obtain the noisy raw LDR images $I^{{}^{\prime}}_{l}$ and $I^{{}^{\prime}}_{s}$, and the corresponding clean raw HDR image $I^{{}^{\prime}}_{gt}$.
+Finally, after the luminance alignment [18], we obtain the noisy raw SDR images $I^{{}^{\prime}}_{l}$ and $I^{{}^{\prime}}_{s}$, and the corresponding clean raw HDR image $I^{{}^{\prime}}_{gt}$.
 Our degraded dataset follows the principle of HDR synthesis—namely, the principle of maximum signal-to-noise ratio.
 In the darkest region, we select long-exposure images; in the brightest region, we select short-exposure images.
-3.2.3 RealRaw-HDR Dataset
+III-C RealRaw-HDR Dataset
 Our dataset is meticulously crafted for dual-exposure HDR sensors, supporting mainstream sensors, including Sony IMX327, IMX385, IMX585, and OV OS05B.
@@
 Fig.
-2 shows an example of two generated LDR-HDR pairs with different exposure ratios.
+2 shows an example of two generated SDR-HDR pairs with different exposure ratios.
 Additionally, by attaching an ISP pipeline to the end of our pipeline, we can create an sRGB-based HDR training dataset.
 In Tab.
-1, we compare the statistics of our dataset with those of other existing HDR datasets.
+I, we compare the statistics of our dataset with those of other existing HDR datasets.
 In this paper, all raw images have been processed with a fixed ISP, and HDR images are processed with the same tone mapping operator to obtain the sRGB version for visualization.
-Table 1: The statistics comparison between Kalanatri [14], Chen [4] and our RealRaw-HDR dataset.
+TABLE I: The statistics comparison between Kalanatri [18], Chen [5] and our RealRaw-HDR dataset.
 Data Quantity Size Format Exposure Ratio
-Kalanatri [14] 74 $1490\times 989$ sRGB 4 &\&& 8 &\&& 16
-Chen [4] 144 $4096\times 2168$ raw, sRGB 4 &\&& 8 &\&& 16
+Kalanatri [18] 74 $1490\times 989$ sRGB 4 &\& 8 &\& 16
+Chen [5] 144 $4096\times 2168$ raw, sRGB 4 &\& 8 &\& 16
 Ours 720 $8192\times 6192$ raw, sRGB 4-16
-3.2.4 Effectiveness of the Data Formation Pipeline
+III-D Effectiveness of the Data Formation Pipeline
 The proposed pipeline for generating HDR data is efficient and user-friendly, allowing easy generation of numerous high-quality data pairs.
 Although our ground truth is derived from a single image, it contains a wide range of information characteristics of HDR images.
-Firstly, the raw images $I_{1}$ and $I_{2}$ are captured by a high dynamic range camera and contain high dynamic range.
-On the other hand, there is a significant difference in the signal-to-noise ratio between HDR images and LDR images.
+Firstly, the raw images $I_{1}$ and $I_{2}$ are captured by an HDR camera.
+On the other hand, there is a significant difference in the signal-to-noise ratio between HDR images and SDR images.
 The exposure-aligned long-exposure images differ from the short-exposure images only in the dark and overexposed regions, in addition to the noise difference.
 Therefore, $I_{l}$ and $I_{s}$ have all the characteristics of real-world long and short-exposure images, and $I_{gt}$ contains a wide range of informative features of real HDR images.
-4 Methodology
-4.1 Overview
+IV Methodology
+IV-A Overview
 HDR reconstruction plays a vital role in various applications, such as mobile photography, high-definition displays, and virtual reality, where lightweight and efficient algorithms are highly demanded due to resource limitations.
@@
 On the other hand, unlike GPU servers, existing optimized manipulations for mobile devices are quite limited, especially on computationally limited devices.
-Unsupported operators have to be processed on the CPU, which not only very low processing speed, but also introduces additional MACs.
-Therefore, we first design a neat UNet with mobile-friendly operations as the base model, then propose a re-parameterizable Topological Convolution Block to improve HDR performance.
+Unsupported operators have to be processed on the CPU, which not only very low processing speed but also introduces additional MACs.
+Therefore, we first design a neat UNet with mobile-friendly operations as the base model, then propose a reparameterizable Topological Convolution Block to improve HDR performance.
 For lightweight design, we do not use the computationally demanding explicit alignment in our HDR network.
 To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss (in Sec.
-4.5) that enables training with unaligned pairs.
-4.2 Base Model
-To ensure high inference speed and cross-device deployment on commodity mobile devices, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a neat UNet consisting of the most basic operations as the base model.
+IV-E) that enables training with unaligned pairs.
+IV-B Base Model
+To pursue high inference speed on commodity mobile devices and the possibility of cross-device deployment, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a neat UNet consisting of the most basic operations as the base model.
 The overall architecture of the base model is shown in Fig.
@@
 In parallel, Encoder-L extracts features from the long-exposure image, offering supplementary features.
+(a) Base Model
+(b) Topology Convolution Block (TCB)
+Figure 3: Illustration of (a) Base Model and (b) Topology Convolution Block (TCB).
+In the training phase, the TCB employs multiple branches, which can be merged into one normal convolution layer in the inference stage.
 Considering the limited bandwidth, we first employ the pixel unshuffle operation [10] to transfer the input raw images $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{s}$ from $C\times H\times W$ to $4C\times\frac{H}{2}\times\frac{W}{2}$ to extract multi-scale contextual information while keeping the MAC of our model as low as possible.
@@
 The network design with low MAC allows for ultra-fast inference on mobile devices, and the basic operation makes cross-device deployment easier.
-(a) Base Model
-(b) Topology Convolution Block (TCB)
-Figure 3: Illustration of (a) Base Model and (b) Topology Convolution Block (TCB).
-In the training phase, the TCB employs multiple branches, which can be merged into one normal convolution layer in the inference stage.
-Figure 4: Visual comparison of state-of-the-art HDR reconstruction methods on our synthetic test dataset.
-4.3 Topological Convolution Block
-Although the plain base model is efficient, its HDR performance is less satisfactory compared to those complicated models, as shown in Tab.
-5.
-We thus employ the re-parameterization technique to enrich the representation capability of the base model.
+IV-C Topological Convolution Block
+Although the plain base model is efficient, its HDR performance is less satisfactory compared to the complicated models, as shown in Tab.
+VIII.
+We thus employ the reparameterization technique to enrich the representation capability of the base model.
 The reparameterization has achieved promising results on other tasks [8, 4, 6, 55].
-We design a flexible re-parameterizable module called the Topological Convolution Block (TCB)), which can more effectively extract edge and texture information for the HDR task.
+We design a flexible reparameterizable module called the Topological Convolution Block (TCB), which can more effectively extract edge and texture information for the HDR task.
 As shown in Fig.
-3(b), the TCB consists of several fundamental units: (1) A standard $3\times 3$ convolution for a solid foundation.
-The standard convolution is denoted as:
+3(b), the TCB consists of several fundamental units:
+(1) A standard $3\times 3$ convolution for a solid foundation. The standard convolution is denoted as:
 $F_{n}=W_{n}*X+B_{n},$ (6)
 where $F_{n}$, $X$, $W_{n}$, and $B_{n}$ represent the output feature, input feature, weights, and bias of the standard convolution, respectively.
-(2) Extending and squeezing convolution to enhance feature expressiveness, the expanding and squeezing feature is extracted as:
+(2) Extending and squeezing the convolution to enhance feature expressiveness, the expanding and squeezing feature is extracted as:
 $F_{es}=W_{s}*(W_{e}*X+B_{e})+B_{s},$ (7)
@@
 (3) Sobel and Laplacian operators for extracting first and second-order spatial derivatives to identify edges, i.e., using a predetermined convolution kernel to process the edge information.
-Denote by $D_{x}$ and $D_{y}$ the horizontal and vertical Sobel filters, $D_{lap}$ is the Laplacian filter.
+Denote by $D_{x}$ and $D_{y}$ the horizontal and vertical Sobel filters, and $D_{lap}$ is the Laplacian filter.
 $D_{x}=\begin{bmatrix}+1&0&-1\\ +2&0&-2\\ +1&0&-1\end{bmatrix}$, (8)
@@
 $F_{j}=X$ (13)
-The output of the TCB in the combination of the five components:
-$F_{TCB}=F_{n}+F_{es}+F_{edge}+F_{c}+F_{j}$ (14)
+The output of the TCB is the combination of the five components:
+$F_{TCB}=F_{n}+F_{es}+F_{edge}+F_{c}+F_{j}.$ (14)
 The combined feature map is then fed into a non-linear activation layer.
@@
 This selective approach is grounded in Laplacian operator effectiveness for noise-free images, underpinning its application to enhance feature representation in contexts devoid of noise.
-4.4 Re-parameterization for Efficient Inference
+IV-D Re-parameterization for Efficient Inference
 To achieve an efficient HDR network that meets the stipulated design prerequisites of low computational complexity and streamlined hardware device deployment, we simplify the TCB reparameterization into a single $3\times 3$ convolution after training.
-Following previous works [6, 5, 44], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolution into a single $3\times 3$ convolution.
+Following previous works [7, 6, 55], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolutions into a single $3\times 3$ convolution.
 Additionally, we combine the Sobel and Laplacian operators into a special $3\times 3$ convolution with a fixed convolution kernel.
@@
 Figure 4: An illustrative sample of data construction for the proposed alignment-free and motion-aware short-exposure-first selection loss.
-Table 2: Performance comparison of different HDR models on our synthetic dataset.
-#Param and FLOPs represent the total number of network parameters and floating-point operations.
-The FLOPs and Run Times results are measured on an RTX 3090 device with a resolution of $4096\times 2952$ raw images.
-Metrics with $\uparrow$ and $\downarrow$ denote higher better and lower better, respectively.
-The best and second-best performances are in bold and underlined, respectively.
-"-" indicates the result is not available.
-Methods FLOPs #Param Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16
-[…]
-DeepHDR[33] 2409.32G 15.26M 4.3 ms 43.3680 1.3767 43.5551 1.3844 43.7312 1.3679 42.8178 1.3779
-NHDRRNet[37] 826.17G 40.26M 7.9 ms 33.0206 2.7308 33.0127 2.7277 33.0392 2.7203 33.0101 2.7443
-UNet-SID[2] 640.89G 7.76M 3.1 ms 43.3892 1.3434 43.3314 1.3535 43.4551 1.3312 43.3811 1.3456
-SGN[8] 712.66G 4.78M 3.3 ms 43.6094 1.3235 43.5074 1.3398 43.7078 1.3067 43.6131 1.3240
-HDR-Transformer[21] 3698.28G 1.23M - 44.3895 1.3681 44.3311 1.3811 44.4140 1.3619 44.4235 1.3613
-AHDRNet[34] 2848.29G 0.93M 23.6 ms 44.7985 1.2939 44.8548 1.2957 44.8343 1.2892 44.7064 1.2968
-Ours 127.55G 0.82M 2.9 ms 44.8081 1.2886 44.7575 1.3000 44.8482 1.2812 44.8187 1.2842
-4.5 Loss Functions
+IV-E Loss Functions
 Alignment-free and motion-aware short-exposure-first selection loss.
@@
 Short-exposure images are often used as reference images in engineering applications.
-Based on the above analysis, we devise the strategies for dual-exposure HDR fusion: Firstly, in scenarios involving motion or overexposure within the fused region, we prefer to select short-exposure image that contains more information; Secondly, when the SNR of the short frame is too low, we prefer to select the long-exposure image that contains more information; Thirdly, our strategy is inclined to address ghost artifacts with a higher priority than lower SNR when ghost artifacts and lower SNR concurrently exist.
+Based on the above analysis, we devise the strategies for dual-exposure HDR fusion: Firstly, in scenarios involving motion or overexposure within the fused region, we prefer to select the short-exposure image that contains more information; Secondly, when the SNR of the short frame is too low, we prefer to select the long-exposure image that contains more information; Thirdly, our strategy is inclined to address ghost artifacts with a higher priority than lower SNR when ghost artifacts and lower SNR concurrently exist.
 For the designed fusion strategy, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate the ghost artifacts.
@@
 Specifically, as shown in Fig.
-5, for each patch of an image, we randomly select a rectangle of random length and width from the long-exposure image and then move and overlap it to a random location in the range of -30 to 30 relative to the patch.
+4, for each patch of an image, we randomly select a rectangle of random length and width from the long-exposure image and then move and overlap it to a random location in the range of -30 to 30 relative to the patch.
 The patch regions before and after the movement are labeled as 1s in the mask $M$.
@@
 $L_{\textrm{pix}}=\|\widetilde{I}^{out}-{I}^{gt}\|_{2},$ (16)
-To achieve the best HDR reconstruction results, we employ the multi-scale structural similarity loss function loss guide model, which learns short-exposure image information for global motion.
+To achieve the best HDR reconstruction results, we employ the multi-scale structural similarity loss function guide model, which learns short-exposure image information for global motion.
 By combining these loss functions, our model effectively produces superior results for both areas with motion and saturated regions.
@@
 We average the two G channels of the output (RGGB pattern) and ground truth (RGGB pattern) respectively, and then concatenate the averaged G channel with the R and B channels to perform a naive transformation to the RGB color space, producing two RGB images: $\widetilde{I}^{out}_{rgb}$ and ${I}^{gt}_{rgb}$.
-Then, we impose the colorfulness loss between the processed output and ground truth by the cosine embedding loss.
+Then, we impose the colorfulness loss between the processed output and the ground truth by the cosine embedding loss.
 $L_{\textrm{b}}=\textrm{Cosine}(\widetilde{I}_{rgb}^{out},{I}_{rgb}^{gt}),$ (18)
@@
 where $\alpha$, $\beta$, $\gamma$, and $\eta$ are the corresponding weight coefficients.
-Figure 6: Visual comparisons with the state-of-the-art methods on HDR sensor raw dataset.
-Table 3: Performance comparison of different HDR models on the actual HDR sensor raw dataset [4]. The best and second-best performances are in bold and underlined, respectively.
+TABLE II: Performance comparison of different HDR models on synthetic dataset from RealRaw-HDR.
+#Param and FLOPs represent the total number of network parameters and floating-point operations.
+The FLOPs and Run Times results are measured on an RTX 3090 device with a resolution of $4096\times 2952$ raw images.
+Metrics with $\uparrow$ and $\downarrow$ denote higher better and lower better, respectively.
+PSNR test in HDR Raw images and $\Delta E$ test in tone-mapped sRGB images.
+The best and second-best performances are in bold and underlined, respectively.
+”-” indicates that inference is not possible due to memory limitations.
+Methods GFLOPs #Param Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16
+[…]
+DeepHDR[44] 2409.32 15.26M 4.3ms 43.3680 1.3767 43.5551 1.3844 43.7312 1.3679 42.8178 1.3779
+NHDRRNet[49] 826.17 40.26M 7.9ms 33.0206 2.7308 33.0127 2.7277 33.0392 2.7203 33.0101 2.7443
+UNet-SID[3] 640.89 7.76M 3.1ms 43.3892 1.3434 43.3314 1.3535 43.4551 1.3312 43.3811 1.3456
+SGN[10] 712.66 4.78M 3.3ms 43.6094 1.3235 43.5074 1.3398 43.7078 1.3067 43.6131 1.3240
+CEN-HDR[36] 162.35 0.19M 2.8ms 42.3021 1.4240 42.5312 1.4237 42.6531 1.4254 41.7219 1.4228
+HDR-GAN[31] 629.09 1.32M 8.3ms 44.7641 1.2972 44.8954 1.2867 44.7743 1.3272 44.6427 1.2778
+HDR-Transformer[25] 3698.28 1.23M - 44.3895 1.3681 44.3311 1.3811 44.4140 1.3619 44.4235 1.3613
+AHDRNet[46] 2848.29 0.93M 23.6ms 44.7985 1.2939 44.8548 1.2957 44.8343 1.2892 44.7064 1.2968
+SCTNet[37] 3025.80 0.97M - 44.8452 1.3048 44.9132 1.3231 44.8712 1.3026 44.7511 1.2889
+Ours 127.55 0.82M 2.9ms 44.8081 1.2886 44.7575 1.3000 44.8482 1.2812 44.8187 1.2842
+Figure 5: Visual comparison of state-of-the-art HDR reconstruction methods on our synthetic dataset from RealRaw-HDR.
+V Experiments
+V-A Experimental Setup
+Datasets and metrics.
+We utilize the proposed RealRaw-HDR dataset for training.
+We first evaluate our method in the synthesized dataset.
+This test set contains 30 samples containing different exposure ratios (i.e., 4, 8, and 16) with a resolution of $4096\times 2176$.
+To validate the validity of our method on real data, we utilize a FUJI-FILM GFX50S II camera to capture seven sets of real-world bracketed exposure raw images and the corresponding static images for generating the ground truth.
+Furthermore, we also utilize the Chen [5] test dataset for cross-validation, which has short- and long-exposure raw pairs captured by a Sony IMX267 image sensor.
+We perform a quantitative evaluation using the PSNR, SSIM, HDR-VDP3 [28], PU [2], ColorVideoVDP [29], and CIE L*a*b* space.
+The HDR-VDP3 [28] metric predicts the quality degradation concerning the reference image.
+We adopt the default settings provided in the HDR-VDP-3 processing framework, which simulate a typical HDR viewing environment.
+Metrics are computed directly on the HDR images without any tone mapping.
+All HDR images are evaluated under a display profile with peak luminance of 1000 cd/m², a contrast ratio of 1000:1, a gamma of 2.2, an ambient illumination of 100 lux, and a screen size of 30 inches at $3840\times 2160$ resolution.
+The viewing distance was set to 0.5 meters.
+To further validate the perceptual quality of the reconstructed HDR images, we incorporated the ColorVideoVDP [29] into our evaluation.
+We used the official implementation (v0.5.0) and applied it to HDR images using the configuration: ”ColorVideoVDP v0.5.0, 75.4 [pix/deg], Lpeak=1000, Lblack=0.5979 [cd/m²], (standard_4k).” CIE L*a*b* space  [^1] CIE L*a*b* is a color space specified by the International Commission on Illumination.
+describes all the colors visible to the human eye and was created to serve as a device-independent model for reference.
+[57, 14, 23] (also known as $\Delta E$).
+It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE L*a*b* color space.
+$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (20)
+where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE L*a*b* version of the predicted HDR image and ground truth, respectively.
+Note that HDR-VDP3, PU, and $\Delta E$ are all tested on sRGB images, which are obtained from HDR Raw images through the same image signal processor process.
+TABLE III: Performance comparison of different HDR models on the actual HDR sensor raw dataset from Chen’s dataset [5].
+PSNR and SSIM tests in HDR Raw images.
+The best and second-best performances are in bold and underlined, respectively.
 Methods All-Exposure Ratio=4 Ratio=8 Ratio=16
@@
 Ours 40.5238 0.9747 1.9568 40.4061 0.9733 1.9743 41.4010 0.9788 1.7974 39.7642 0.9721 2.0988
-5 Experiments
-5.1 Experimental Setup
-Datasets and metrics.
-We utilize the proposed RealRaw-HDR dataset for training.
-We first evaluate our method in the synthesized dataset.
-This test set contains 30 samples containing different exposure ratios (i.e., 4, 8, and 16) with a resolution of $4096\times 2176$.
-To validate the validity of our method on real data, we utilize a FUJI-FILM GFX50S II camera to capture seven sets of real-world bracketed exposure raw images and the corresponding static images for generating the ground truth.
-Furthermore, we also utilize the Chen [4] test dataset for cross-validation, which has short- and long-exposure raw pairs captured by a Sony IMX267 image sensor.
-We perform a quantitative evaluation using the PSNR, SSIM, and CIE L*a*b* space  [^1] CIE L*a*b* is a color space specified by the International Commission on Illumination.
-It describes all the colors visible to the human eye and was created to serve as a device-independent model for reference.
-[43, 11, 19] (also known as $\Delta E$).
-$\Delta E$ can effectively assess chromaticity, contrast, and color accuracy variations within HDR images rather than exclusively concentrating on luminance differences (HDR-VDP-2).
-It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE L*a*b* color space.
-$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (20)
-where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE L*a*b* version of the predicted HDR image and ground truth, respectively.
+TABLE IV: Performance comparison of different HDR models on the actual HDR sensor raw dataset from Chen’s dataset [5].
+All HDR images are now evaluated under a 1000 $cd/m^{2}$ display profile, with metrics computed on the HDR image without tone mapping.
+The best and second-best performances are in bold and underlined, respectively.
+Methods All-Exposure Ratio=4 Ratio=8 Ratio=16
+[…]
+DeepHDR[44] 9.034 38.180 9.646 9.042 37.581 8.939 39.136 9.121 37.824
+NHDRRNet[49] 8.771 30.743 9.276 8.687 30.756 8.739 30.946 8.885 30.567
+UNet-SID[3] 9.089 38.083 9.701 9.087 37.613 8.986 38.969 9.194 37.665
+SGN[10] 9.104 37.696 9.705 9.074 37.376 9.010 38.368 9.230 37.344
+HDR-Transformer[25] 9.106 38.145 9.697 9.093 37.694 9.028 38.752 9.197 37.988
+AHDRNet[46] 9.153 38.220 9.734 9.189 37.876 9.037 38.756 9.234 38.038
+Ours 9.138 39.276 9.702 9.108 39.459 9.060 39.919 9.247 38.449
+Figure 6: Visual comparisons with the state-of-the-art methods on the actual HDR sensor raw dataset from Chen’s dataset [5].
 Implementation details.
@@
 The input patch size for the network is $256\times 256$, and the batch size is 32.
-Our model is implemented in PyTorch and trained with an NVIDIA RTX 3090 GPU.
-5.2 Comparison with the Other Methods
-We choose several representative low-level vision methods for comparisons, including four HDR methods based on sRGB images (AHDRNet [34], DeepHDR [33], NHDRRNet [37], HDR-Transformer[21]), as well as two methods for denoising raw images (SGN [8] and UNet-SID [2]).
-For fair comparisons, we re-train all the methods using the RealRaw-HDR dataset.
-Additionally, for AHDRNet, DeepHDR, HDR-Transformer, and NHDRRNet, we modify the network inputs to accommodate dual-exposure raw images.
+Our model is implemented in PyTorch and trained with an NVIDIA RTX 3090 GPU with Intel Xeon Platinum 8369B CPU (64 vCPUs@2.90GHz).
+The total training time is 60 hours.
+V-B Comparison with the Other Methods
+We choose several representative low-level vision methods for comparisons, including six HDR methods based on sRGB images (AHDRNet [46], DeepHDR [44], NHDRRNet [49], HDR-GAN [31], HDR-Transformer [25], CEN-HDR [36], and SCTNet [37]), as well as two methods for denoising raw images (SGN [10] and UNet-SID [3]).
+For fair comparisons, we retrain all the methods using the RealRaw-HDR dataset.
+Additionally, for AHDRNet, DeepHDR, HDR-GAN, HDR-Transformer, and NHDRRNet, we modify the network inputs to accommodate dual-exposure raw images.
 Similarly, for SGN and UNet-SID, we concatenate the long- and short-exposure raw pairs as inputs.
-Evaluation on synthetic dataset.
+Evaluation on the synthetic dataset.
 We first evaluated our method on a synthetic dataset generated using the raw HDR data formation pipeline.
 The quantitative comparison results are shown in Tab.
-2.
-The results clearly show that our method outperforms previous methods in almost all metrics on the synthetic dataset.
-Notably, our lightweight-efficient RepUNet model has fewer parameters (0.82M) and remarkably low GFLOPs (127G FLOPs).
-This efficiency allows us to process two 4K Bayer raw images at only 2.9ms using an NVIDIA RTX 3090 GPU.
-In contrast, other models with comparable performance necessitate significantly longer processing times.
-RepUNet achieves comparable performance with AHDRNet with only $4.5\%$ of its computational complexity (127G vs.
-2848G).
-Fig 4 visualizes that our method can effectively eliminate noise and ghosting artifacts in the reconstructed HDR.
+II.
+The results demonstrate that our method achieves performance comparable to the state-of-the-art (SOTA) method on the synthetic dataset.
+Notably, our lightweight and efficient RepUNet architecture significantly reduces both model size and computational complexity, with only 0.82M parameters and 127 GFLOPs.
+This high efficiency enables the processing of 4K Bayer raw images in just 2.9 ms on an NVIDIA RTX 3090 GPU, outperforming other models with similar performance that require substantially longer processing times.
+Compared to CEN-HDR, RepUNet improves PSNR by 2.5 dB while maintaining comparable computational costs.
+Furthermore, RepUNet achieves performance on par with AHDRNet, SCTNet, and HDR-GAN, but with dramatically lower computational complexity: it requires only 4.5% of the computational effort of AHDRNet (127 GFLOPs vs.
+2848 GFLOPs), 4.2% of SCTNet (127 GFLOPs vs.
+2025 GFLOPs), and 20.2% of HDR-GAN (127 GFLOPs vs.
+629 GFLOPs).
+These results highlight the superior efficiency and practicality of our approach.
+Fig.
+5 shows that our method can effectively eliminate noise and ghosting artifacts in the reconstructed HDR.
 In comparison, DeepHDR [44], NHDRRNet [49], and SGN [10] exhibit numerous artifacts in the palm motion region.
-However, our proposed RepUNet can reconstruct HDR images without ghosting (see rows 2 in Fig.
-4).
-Figure 7: Visual comparison of state-of-the-art HDR reconstruction methods on FUJI raw dataset.
-Table 4: Quantitative comparisons of different loss functions. AMSS-Loss represents the alignment-free and motion-aware short-exposure-first selection loss.
-ID Method Bayer-Loss AMSS-Loss All-Exposure Ratio=4 Ratio=8 Ratio=16
-[…]
-1 RepUNet 39.6403 2.1479 39.5214 2.1577 40.5216 1.9603 38.8780 2.3256
-2 RepUNet 40.0251 2.1162 39.9900 2.1143 40.8125 1.9700 39.2729 2.2642
-3 RepUNet 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
-Table 5: Reparameterization ablation results. The FLOPs and run times are measured on the raw image with a 4K resolution.
-Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
-[…]
-Base Model 93.26G 0.82M 3.0 ms 39.7941 2.1202 39.8026 2.1036 40.5092 1.9790 39.0705 2.2780
-RepUNet 93.26G 0.82M 2.9 ms 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
+However, our proposed RepUNet can reconstruct HDR images without ghosting (see row 2 in Fig.
+5).
+Figure 7: Visual comparison of state-of-the-art HDR reconstruction methods on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.
 Evaluation on HDR sensor dataset.
-To validate the validity of our method on the real-world HDR sensor dataset, we utilize the Chen [4] test dataset for cross-validation, which has raw images captured by a Sony IMX267 image sensor.
-Compared with previous methods, our method achieves state-of-the-art performance in visual quality and quantitative metrics.
+To validate the validity of our method on the real-world HDR sensor dataset, we utilize the Chen test dataset [5] for cross-validation, which has raw images captured by a Sony IMX267 image sensor.
+Our method achieves comparable performance in visual quality and quantitative metrics compared to previous methods.
 The visual results from tests on the HDR sensor raw dataset (as shown in Fig.
-6) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from severe color casts.
+6) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from color casts.
 Furthermore, results in Tab.
-3 reveal that compared to AHDRNet [34], our method yields an improvement of more than 0.35 dB and 0.05 gain in PSNR and $\Delta E$, respectively, for scenes with an exposure ratio of 8.
+III reveal that compared to AHDRNet [46], our method yields an improvement of more than 0.35 dB and 0.05 gain in PSNR and $\Delta E$, respectively, for scenes with an exposure ratio of 8.
 On average, our method attains gains exceeding 0.1 dB and 0.15 in PSNR and $\Delta E$.
-Table 6: Performance comparison of different HDR models on RealRaw-HDR dataset with realistic exposure ratios.
+In addition, as shown in Tab.
+IV, RepUNet far outperforms the other methods in terms of PU-PSNR metrics, outperforming the suboptimal method by 1 dB.
+While our method yields a slightly lower score on ColorVideoVDP, it consistently outperforms competing methods on other HDR metrics (e.g., PSNR, SSIM, LPIPS), and offers lower computational cost and faster inference.
+TABLE V: Performance comparison of different HDR models on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.
 The FLOPs are measured on the raw image of $7808\times 5824$ resolution.
@@
 We then evaluate our method on the FUJI raw datasets, which are real-world bracketed exposure raw images captured by the FUJI-FILM GFX50S II camera.
+Compared with previous methods, our method achieves state-of-the-art performance in visual quality and quantitative metrics.
 Fig.
-7 compares results from two high dynamic range scenes, where our method achieves significantly better visualization.
+7 compares results from HDR scenes, where our method achieves significantly better visualization.
 Our method can recover both fine details in overexposed regions and rich colors in underexposed areas without introducing artifacts (see rows 1 and 2).
 Compared to AHDRNet, our method can effectively remove noise and preserve the structure of dark regions.
+Tab.
+V shows that our method is far less computationally (FLOPs) intensive than AHDRNet [46] and HDR-Transformer at similar PSNR.
 Notably, the alignment module in DeepHDR, AHDRNet, HDR-Transformer, and NHDRRNet requires many line buffers, making it challenging to deploy on resource-limited edge devices.
 HDR-Transformer fails to perform inference even on RTX 3090 devices.
-In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images, as evident in the visual results of Fig.
-4, Fig.
-6, and Fig.
-7.
+In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images.
 Although all models are trained on the RealRaw-HDR dataset, which is synthesized using the data formation pipeline, they consistently excel on both the synthetic test dataset and the real-world dataset.
@@
 These results are solid evidence of the generalizability of our proposed RealRaw-HDR dataset and the HDR data formation pipeline.
-5.3 Ablation Study
-This section investigates the raw LDR-HDR pair formation pipeline and the importance of different components in the whole RepUNet.
+V-C Ablation Study
+This section investigates the raw SDR-HDR pair formation pipeline and the importance of different components in the whole RepUNet.
 We ablate the baseline model step by step and compare the performance differences.
-Generalization of our LDR-HDR pair formation pipeline.
-Our raw LDR-HDR pair formation pipeline is proposed to generate paired raw LDR-HDR data but can also be adapted to generate paired sRGB HDR data.
+Generalization of our SDR-HDR pair formation pipeline.
+Our raw SDR-HDR pair formation pipeline is proposed to generate paired raw SDR-HDR data, but can also be adapted to generate paired sRGB HDR data.
 To demonstrate such generalization, we transform the collected RealRaw-HDR dataset with a fixed ISP pipeline into the sRGB color space, named the Raw2RGB-HDR dataset.
@@
 Results in Tab.
-7 show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on the Kalantari dataset by 2.86 dB in PSNR.
+VI show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on the Kalantari dataset by 2.86 dB in PSNR.
 The performance gains benefit from an efficient and user-friendly data acquisition pipeline that generates more trainable data pairs.
-The results demonstrate that our data pipeline is also effective in generating paired LDR-HDR data in sRGB space.
-Table 7: We train the sRGB HDR method AHDRNet on our Raw2RGB-HDR dataset and Kalantari dataset, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR dataset.
+The results demonstrate that our data pipeline is also effective in generating paired SDR-HDR data in sRGB space.
+TABLE VI: We train the HDR sRGB method AHDRNet on our Raw2RGB-HDR and Kalantari datasets, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR.
 Method Dataset PSNR PSNR-$\mu$
@@
 Raw2RGB-HDR 38.3183 39.8896
+TABLE VII: Quantitative comparisons of different loss functions. AMSS-Loss represents the alignment-free and motion-aware short-exposure-first selection loss.
+ID Method Bayer-Loss AMSS-Loss All-Exposure Ratio=4 Ratio=8 Ratio=16
+[…]
+1 RepUNet 39.6403 2.1479 39.5214 2.1577 40.5216 1.9603 38.8780 2.3256
+2 RepUNet 40.0251 2.1162 39.9900 2.1143 40.8125 1.9700 39.2729 2.2642
+3 RepUNet 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
 Figure 8: Visual results of RepUNet and its baseline variants. Combining these loss functions allows our model to produce top-notch results for motion and saturated areas effectively.
+TABLE VIII: Reparameterization ablation results. The FLOPs and run times are measured on the raw image with a 4K resolution.
+Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
+[…]
+Base Model 93.26G 0.82M 3.0 ms 39.7941 2.1202 39.8026 2.1036 40.5092 1.9790 39.0705 2.2780
+RepUNet 93.26G 0.82M 2.9 ms 40.5238 1.9568 40.4061 1.9743 41.4010 1.7974 39.7642 2.0988
 Loss functions.
-To test the effects of alignment-free and motion-aware short-exposure-first selection loss and Bayer loss, we set the $L2$ joint $L_{\textrm{ssim}}$ loss as the baseline loss and step-by-step modify the loss function combination.
+To test the effects of alignment-free and motion-aware short-exposure-first selection loss and Bayer loss, we set the $L2$ joint $L_{\textrm{ssim}}$ loss as the baseline and step-by-step modify the loss function combination.
 Tab.
-4 and Fig.
+VII and Fig.
 8 show that adding the AMSS and Bayer loss steadily improves visual quality and quantitative results.
@@
 […]
-6 Conclusion
-In the paper, we proposed a Topological Convolution Block (TCB) for efficient and light-weight HDR design for mobile devices.
+VI Conclusion
+In the paper, we proposed a Topological Convolution Block (TCB) for an efficient and lightweight HDR design that may be suitable for mobile devices.
 Based on the proposed TCB, we further designed RepUNet, aiming at balancing hardware efficiency and PSNR/SSIM indexes.
-Furthermore, We propose a novel computational photography based pipeline for raw HDR image formation and construct a real-world raw HDR dataset – RealRaw-HDR.
-Meanwhile, we designed plug-and-play alignment-free and motion-aware short-exposure-first selection loss efficient mitigate ghost artifacts.
-Our empirical evaluation validates the effectiveness of the proposed LDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.
+Furthermore, we propose a novel computational photography-based pipeline for raw HDR image formation and construct a real-world raw HDR dataset – RealRaw-HDR.
+Meanwhile, we designed a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate ghost artifacts.
+Our empirical evaluation validates the effectiveness of the proposed SDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.
+VII Limitations and Discussion
+In this work, we propose an efficient and lightweight HDR network for dual-exposure HDR sensors.
+We have designed the TCB to prioritize computational efficiency, but real-world mobile deployments require additional hardware-specific adaptations.
+We will address this issue in future work through real-device testing.
+In addition, the proposed data synthesis pipeline only supports dual-exposure HDR sensor settings, and we will explore data synthesis pipelines for a wider range of triple-exposure HDR sensors in subsequent research.
+Additionally, we will focus on exploring more complex application scenarios such as flickering lights and large motion scenes.
+Appendix
+In this Appendix, we provide additional results and analysis.
+VIII Supplementary Materials
+Fig.
+9 to Fig.
+12 show additional visual comparisons of our method with state-of-the-art HDR reconstruction methods.
+As shown in these figures, DeepHDR, NHDRRNet, and SGN all have obvious ghosting and color residuals, especially for moving color objects, and the reconstruction results will have severe ghosting.
+Our proposed RepUNet can reconstruct HDR images without ghosting and color bias, and we can effectively remove the noise in the reconstructed HDR.
+Figure 9: Visual comparisons with the state-of-the-art methods on the actual HDR sensor raw dataset from Chen’s dataset [5].
+Figure 10: Visual comparison of state-of-the-art HDR reconstruction methods on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.
+Figure 11: Visual comparison of state-of-the-art HDR reconstruction methods on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.
+Figure 12: Visual comparison of state-of-the-art HDR reconstruction methods on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.